### v3

# Nash-equilibrium Seeking Algorithm for Power-Allocation Games on Networks of International Relations

Abstract
In the field of international security, understanding the strategic interactions between countries within a networked context is crucial. Our previous research has introduced a “games-on-signed graphs [LiMorse2022]” framework to analyze these interactions. While the framework is intended to be basic and general, there is much left to be explored, particularly in capturing the complexity of strategic scenarios in international relations. Our paper aims to fill this gap in two key ways. First, we modify the existing preference axioms to allow for a more nuanced understanding of how countries pursue self-survival, defense of allies, and offense toward adversaries. Second, we introduce a novel algorithm that proves the existence of a pure strategy Nash equilibrium for these revised games. To validate our model, we employ historical data from the year 1940 as the game input and predict countries’ survivability. Our contributions thus extend the real-world applicability of the original framework, offering a more comprehensive view of strategic interactions in a networked security environment.
{IEEEkeywords}
Signed network, network games, Nash equilibrium, network dynamics
1 Introduction
1.1 Background and motivation
The study of international relations has long been a dynamic and evolving field of inquiry. In recent years, the growing complexity and volatility of global interactions have prompted increased interest in quantitative approaches aimed at objectively characterizing the underlying mechanisms of international relations. This methodological shift toward empirical and model-based analysis has emerged as a prominent research trend, with scholars striving to develop frameworks capable of capturing the dynamic and strategic nature of interstate interactions[deutsch1964multipolar, cranmerKantianFractionalizationPredicts2015]. Within this context, complex network theory serves as a vital tool for describing international relations. On the one hand, it enables the representation of relationships among diverse actors, thereby uncovering the fundamental macro-structures underlying global politics. On the other hand, it provides a basis for quantitative decision-making in international contexts. By utilizing the complex network framework, researchers can abstract away from some of the intricacies of real-world international relations, concentrate on core strategic dynamics, and simultaneously incorporate multiple influencing factors into their analysis—thus progressively aligning the model with real-world phenomena. International relations can be understood as a complex network of interactions among states, shaped by evolving alliances, conflicts, and strategic exchanges. Within this framework, each state aims to maximize its own payoff. Game-theoretic approaches offer a particularly well-suited analytical framework for examining such behavior, providing valuable insights into the strategic actions of states operating within this networked environment.
The application of game theory to international relations has thus attracted increasing scholarly attention, providing robust analytical frameworks for modeling strategic decision-making and anticipating outcomes in an ever more interconnected global landscape. As international interactions become increasingly complex, game-theoretic approaches have not only gained prominence but have also become indispensable for deepening our understanding of geopolitical dynamics and guiding effective policy development. Among the various challenges addressed in this context, the strategic allocation of limited resources stands out as particularly pressing—driven by the intensifying global competition for influence and the deepening interdependence among countries.
Despite the growing popularity of using complex networks and game theory to study issues in international relations, existing research still faces several limitations that need to be addressed. Some studies analyze the changes in international relations from a qualitative perspective[goddardBrokeringChangeNetworks2009, hafner-burtonNetworkAnalysisInternational2009]. For analytical tractability, they deemphasize the strategic behavior of individual nodes, with more attention paid to the network’s overall characteristics and topology.
To better capture the complexity of strategic scenarios in international relations, first we refine the framework of the power-allocation game based on the work of Li and Morse[LiMorse2017_CDC, LiMorse2018_AutomaticaSinica, LiMorse2022]. In our model, we respectively represent countries and their relationships using a signed graph composed of nodes and edges. Each country holds a non-negative quantity of power, which can be allocated either to support friendly countries or to attack antagonistic countries. The way power is allocated forms the strategy of each country. And the strategies chosen by all countries determine each country’s state, which can be classified as safe, precarious, or dangerous. A country which is not dangerous prefers to maximize the sum of its friendly countries that are not dangerous and its antagonistic countries that are not safe. We prove the existence of the Nash Equilibrium of this model and proposed an algorithm to construct a special kind of Nash Equilibria. Moreover, we generalize the original model into a dynamic setting where countries can update their power and allocation strategies. A country’s power increases if it is safe, and decreases if it is dangerous. We establish the convergence of the proposed dynamic model. In the final part of this paper, we carried out simulations on the two models to examine their effectiveness in real-world scenarios.
1.2 Literature review
In recent years, there has been growing interest in applying mathematical modeling as a research approach in the field of international relations. The international system can be viewed as a network of interactions among countries, organizations, and other global actors. These systems are inherently dynamic, influenced by factors such as national power, political ideologies, and shifting alliances, all of which influence cooperation patterns, power structures, and conflict resolution mechanisms. Foundational work in this area was carried out by scholars such as Karl Deutsch, whose seminal work[deutsch1963nerves, deutsch1968toward, deutsch1975some] advocated for the use of formal modeling in international relations [bryen2012application]. For example, he employed formal, semi-quantitative models to explore a long-standing yet empirically underexamined question in the field: whether multipolar systems are more stable than bipolar ones[deutsch1964multipolar].
Many core research questions in international relations can, at their foundation, be framed within a game-theoretic framework. Countries may be viewed as strategic players who must decide how to allocate limited resources—whether to support friendly countries or attack hostile ones—while striving to improve their overall position in the international system. These strategic objectives correspond closely with the concept of player preferences in game theory.
However, existing models frequently fall short in capturing the full complexity of these interactions. Some of them focus only on games involving two or three players or two or three strategies[kyddWhichSideAre2003, correaGameTheoryInstrument]. Even though some extensions to more general situations are mentioned, there is a lack of formal game-theoretic structure[decanioGameTheoryClimate2013, jervisRealismGameTheory1988], and others overlook the intricate network of relationships that exist among countries[snidalGameTheoryInternational1985]. These shortcomings highlight the need for more general and realistic models capable of representing the deeper dynamics that drive strategic behavior in international relations.
In today’s world, marked by growing resource demands and increasing complexity in international relations, understanding strategic resource allocation has become essential for policymakers, scholars, and practitioners alike. The exercise of national power now takes place in an environment marked by mutual vulnerabilities, necessitating careful calculation, prudence, and strategic restraint. This perspective aligns with the literature on network games (e.g.,[BjornerLovaszShor1991, BjornerLovasz1992, ChaudhuriGrahamJamall2008, KalaiZemel1982, Myerson1977, DubeyShapley1984, PachecoTraulsenNowak2006]), where resource allocation over networks has been extensively examined. These studies span a wide range of domains, including transportation systems and wireless communication networks (e.g.,[BredinMaheswaranImerBasarKotzRus2000, EtesamiBasar2014, JohariTsitsiklis2004, CominettiCorreaStierMoses2006]).
However, these existing models are often not directly applicable to the types of problems encountered in international politics. On the one hand, many of these models (e.g.,[myersonGraphsCooperationGamesa, harksEquilibriumComputationResource2022]) are built on the assumption of purely competitive relationships among agents, overlooking the complex combination of cooperation and conflict that defines international interactions. On the other hand, most of them (e.g.,[nguyenPriceBasedResourceAllocation2021]) define payoffs solely from a resource-distribution perspective, where an agent’s total payoff is simply the sum of benefits received from each resource. This formulation neglects how the inherent characteristics of the agents themselves influence their payoffs.
In 2018, Li and Morse proposed a “games-on-signed-graphs” framework[LiMorse2022], which models how countries strategically allocate resources—primarily national power—to achieve favorable outcomes in an increasingly complex international environment. Within this framework, they introduced a preference structure under which a country adjusts its strategy only if doing so increases its total gain without reducing its benefit from any other country.
Building on this foundation, we propose an enhanced preference model to more closely reflect the observed behavior and strategic priorities of countries in real-world contexts. Inspired by the notion of influence maximization, our revised static game formulation allows countries not only to pursue self-survival but also to maximize the number of safe friendly countries and the number of dangerous antagonistic countries. This addition introduces a more comprehensive and strategic logic to country behavior within the game.
Furthermore, by introducing a dynamic adjustment mechanism that allows countries to update their strategies over time, we extend the static model into a dynamic framework. This transformation enhances the model’s applicability and relevance to real-world international dynamics, enabling it to capture temporal patterns in strategic decision-making.
1.3 State of contribution
In this paper, we conduct a comprehensive analysis of the two models we proposed in order to prove the existing of the Nash equilibrium of the static model and the convergence of the dynamic model. In addition we simulate to validate the models’ ability to predict real-world outcomes. The specific contributions of this paper include the following aspects.
Firstly, we extended the power-allocation game model to broaden its applicability. Compared with Li’s work, our revised model allow countries to forgo benefits from some friendly countries in exchange for greater individual gains. This adjustment introduces a wider range of feasible strategies for each country. Furthermore, we generalize the static model into a dynamic framework in which a country’s power evolves in response to its current state. This dynamic approach captures the temporal development of strategic interactions, providing a more realistic depiction of how countries adapt their behavior over time. Such a framework more accurately reflects real-world political and strategic environments. In addition, the dynamic model facilitates the analysis of long-term trends, stability, and convergence properties, thereby improving the model’s relevance to practical applications and policy evaluation.
Secondly, we conduct a thorough theoretical analysis for both models. For the static model, we prove the existence of a Nash equilibrium, highlighting its theoretical relevance and offering a stable foundation for further investigation. In addition, we propose a novel algorithm to prove the existence of pure strategy Nash equilibrium for this new game. The resulting Nash equilibrium illustrates that all countries can remain safe even without support from friendly countries. Moreover, we prove that the dynamic process will almost surely converge to a steady state in terms of any countries’ power and strategy matrix, providing strong support for the model’s theoretical robustness and practical utility. Notably, the equilibrium reached by the dynamic model corresponds to the Nash equilibrium of the static model with the same post-convergence conditions.
Thirdly, to support the application of the static model in real-world scenarios, we introduce the concept of a country’s survival likelihood, defined as the proportion of the Nash equilibria where the country is non-dangerous to the total Nash equilibria. Given the international relationships and national power in a specific year, we conduct simulations to estimate the approximate survival likelihood for each country. Simultaneously, we assess whether a country is dangerous based on historical data and a unified assessment criterion. By comparing the simulation results with historical facts, we obtain a measure of the model’s predictive accuracy. During the World War II period (1939–1945), our model achieved an accuracy rate exceeding $70\%$.
For the dynamic model, we apply survival likelihood to predict changes in GDP, thereby evaluating the validity of the model’s representation of power dynamics. To quantify the accuracy, we calculate the percentage of countries that are both safe and show GDP growth, as well as those that are dangerous and show GDP decline. From 1954 to 2009, the accuracy consistently remained above $70\%$, with most years exceeding $80\%$, demonstrating the reliability of using survival likelihood to predict changes in national power. Additionally, we simulate to investigate how the average power varies with the network’s edge density, exploring what types of networks can promote overall prosperity among all countries. These findings underscore the potential of our models as valuable tools for analyzing international relations and informing policy decisions.
1.4 Organization
The remainder of this paper is organized as follows: Section II introduces the formal definitions and setup of the power-allocation game model, including the existence of pure strategy Nash equilibria, a method for computing such equilibria, and an analysis of their properties under specific network structures. Section III extends the model by formulating a dynamic process and providing the corresponding theoretical analysis. Section IV presents simulation results for both the static and dynamic models. Finally, Section V concludes the paper.
2 Static model: power-allocation games on signed networks
2.1 Notations, definitions and model setup
Denote by $\mathbb{R}_{\geq 0}$ the set of non-negative real numbers. Denote by $\mathbb{N}_{+}$ the set of positive integer and let $\mathbb{N}=\mathbb{N}_{+}\cup\{0\}$. Denote by $\text{card}(\cdot)$ the number of elements in a set. Let $\mathbf{1}_{\{\cdot\}}$ be the indicator function, i.e., $\mathbf{1}_{\{C\}}=1$ if the condition $C$ holds, and $\mathbf{1}_{\{C\}}=0$ otherwise. Denote by $\text{diag}(x)$ the diagonal matrix with the entries of the vector $x$ on its diagonals. Define the sign function
[…]
In this section, we formalize our power-allocation game on signed networks in the context of international relations.
International relations as a signed graph: Consider a world of $n$ countries, indexed by $i\in\mathcal{V}=\{1,\dots,n\}$, whose mutual relations are fixed. For each country $i$, let $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ denote its sets of friendly and antagonistic countries, respectively. All relations are assumed bilateral and symmetric: $j\in\mathcal{F}_{i}$ if and only if $i\in\mathcal{F}_{j}$, and $j\in\mathcal{A}_{i}$ if and only if $i\in\mathcal{A}_{j}$. For convenience, we also let $i\in\mathcal{F}_{i}$ for all $i\in\mathcal{V}$. The $n$ countries together with the sets $\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ and $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$ form an undirected and unweighted signed graph, where nodes represent countries and positive (negative) edges represent friendly (antagonistic) relations. The graph is not assumed to be complete, i.e., $\mathcal{A}_{i}\cup\mathcal{F}_{i}=\mathcal{V}$ does NOT necessarily hold for every $i$. Throughout this paper, we use the terms “graph” and “network” (as well as “node” and “country”) interchangeably.
Power-allocation strategies: For each country $i$, let its power be a non-negative value $p_{i}$, which is assumed fixed for now. In the next section, we will model how $p_{i}$ co-evolves with the countries’ strategies. Country $i$ allocates its power either to attack its antagonistic countries or to support its friendly ones. For any $j\in\mathcal{A}_{i}$ (or $j\in\mathcal{F}_{i}$, respectively), let $x_{ij}\geq 0$ denote the amount of power that country $i$ spends on attacking (or supporting) $j$. For simplicity, both $p_{i}$ and $x_{ij}$ are assumed to take values from $\mathbb{N}$ for all $i,j\in\mathcal{V}$. This assumption of quantized power allocation does not alter the essence of the model and sometimes can even be more realistic, since national power, such as human resources or military units, is not always continuously divisible. The row vector $(x_{i1},\dots,x_{in})$ is referred to as the strategy of country $i$. Denote by $X=(x_{ij})_{n\times n}$ the strategies of all countries and call it a strategy matrix. By definition,
$X\in\Omega=\Big\{X\in\mathbb{N}^{n\times n}\,\Big|\,$ $\text{For any }i,\ \sum_{j\in\mathcal{A}_{i}\cup\mathcal{F}_{i}}x_{ij}=p_{i}$
$\text{and }x_{ij}=0\text{ if }j\notin\mathcal{A}_{i}\cup\mathcal{F}_{i}\Big\}.$
States of countries: Given a strategy matrix $X$, each country’s state is determined as follows.
Definition 1
(Countries’ States) Given any strategy matrix $X\in\mathbb{N}^{n\times n}$, for any $i\in\mathcal{V}$, define
$s_{i}(X)=\sum\limits_{j\in\mathcal{A}_{i}}x_{ij}+\sum\limits_{j\in\mathcal{F}_{i}}x_{ji}-\sum\limits_{j\in\mathcal{A}_{i}}x_{ji}.$
The state of each country $i$ is either safe, precarious, or dangerous, determined as follows:
$\text{Country}\ i\ \text{is}\begin{cases}\text{safe},&\text{if}\ \ s_{i}(X)>0,\\ \text{precarious},&\text{if}\ \ s_{i}(X)=0,\\ \text{dangerous},&\text{if}\ \ s_{i}(X)<0.\end{cases}$
To put it simply, the state of a country depends on the power it spends on attacking its antagonistic countries, plus the support from its friendly countries, and minus the attack from its antagonistic countries.
Preference axioms: To make our model more general and inclusive, we do not specify the countries’ utility functions, but instead assume they obey the following preference axiom.
Definition 2
(Preference) Given two strategy matrices $X$ and $Y$, country $i\in\{1,\dots,n\}$ prefers $X$ to $Y$, denoted by $X\geq_{i}Y$, if either of the following two conditions hold:
  1. 1.
    $s_{i}(X)\geq 0$ and $s_{i}(Y)<0$;
  2. 2.
    Either $s_{i}(X),s_{i}(Y)\geq 0$ or $s_{i}(X),s_{i}(Y)<0$. Moreover,
    $\quad\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{s_{j}(X)\leq 0\}}+\sum_{l\in\mathcal{F}_{i}}\mathbf{1}_{\{s_{l}(X)\geq 0\}}$
    $\geq\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{s_{j}(Y)\leq 0\}}+\sum_{l\in\mathcal{F}_{i}}\mathbf{1}_{\{s_{l}(Y)\geq 0\}}.$
Remark 1
There exist various utility functions compatible with the axioms above, e.g., the utility $u_{i}(X)=n\mathbf{1}_{\{s_{i}(X)\geq 0\}}+\sum_{j\in\mathcal{F}_{i}}\mathbf{1}_{\{s_{j}(X)\geq 0\}}+\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{s_{j}(X)\leq 0\}}$ for each $i\in\mathcal{V}$.
Compared with the power-allocation game proposed in [LiMorse2022], the static game analyzed in this paper introduces a key conceptual advancement: countries are allowed to act in a more strategic and realistic manner. In [LiMorse2022], each country would revise its decision only if it could increase its total payoff without reducing its benefit from any other country. In contrast, our model relaxes this overly conservative assumption by introducing a more flexible preference axiom: one that permits a country to sacrifice the safety of certain friendly states when doing so leads to greater overall self-interest. This modification captures the essence of real-world strategic behavior, where national decisions often involve balancing cooperative commitments with self-serving objectives.
Nash equilibrium: By Definition 2, a strategy matrix $X^{*}\in\Omega$ is called a pure-strategy Nash equilibrium if, for any $i\in\mathcal{V}$, $X^{*}\geq_{i}X$ for any $X\in\Omega$ that that may differ from $X^{*}$ only in the $i$-th row.
2.2 Existence and computation of Nash equilibrium
In this section, we establish one of the main results of this paper. That is, any game satisfying the preference axioms given by Definition 2 admits at least one Nash equilibrium. The theorem is formally presented as follows.
Theorem 1 (Existence of Nash Equilibrium)
Given $n$ countries with powers $p_{1},\dots,p_{n}$ and their respective sets of friendly and antagonistic countries $\mathcal{F}_{1},\mathcal{A}_{1},\dots,\mathcal{F}_{n},\mathcal{A}_{n}$, the static power-allocation game defined in Section 2.1 admits at least one pure-strategy Nash equilibrium. In this equilibrium, no country is dangerous, and no power is allocated between any pair of friendly countries.
The proof proceeds in three steps, outlined as follows:
  1. 1.
    Construct a subset of strategy matrices, termed the no-support-no-unsafe (NSNU) matrices.
  2. 2.
    Define an iterative procedure from an NSNU matrix and based on an operation called the preferable adjustment.
  3. 3.
    Show that this iteration must terminate in finite time, and that the resulting strategy matrix constitutes a Nash equilibrium of the static power-allocation game.
Following the above outline, some useful definitions and facts are introduced first.
Definition 3 (NSND strategy matrix)
Consider the static power-allocation game defined in Section 2.1. A strategy matrix $X\in\Omega$ is called a no-support-no-dangerous (NSND) strategy matrix, if it satisfies that 1) $X=X^{\top}$; 2) $x_{ij}=0$ for any i,ji,\,j such that $j\notin\mathcal{A}_{i}\cup\{i\}$.
The following facts can be easily inferred from the definition of countries’ states and the preference axioms.
Fact 1
If $X$ is an NSND strategy matrix, then $x_{ij}=0$ for any $i\in\mathcal{V}$ and any $j\in\mathcal{F}_{i}\setminus\{i\}$. As a result, with $X=X^{\top}$, any country $i\in\mathcal{V}$ is either safe (when $x_{ii}>0$) or precarious (when $x_{ii}=0$). That is, no country is dangerous under any NSND strategy matrix.
Fact 2
For any NSND strategy matrices $X$ and $Y$, for any $i\in\{1,\dots,n\}$, $X\geq_{i}Y$ if and only if
$\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}=0\}}$ $=\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{s_{j}(X)=0\}}$
$\geq\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{s_{j}(Y)=0\}}=\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{y_{jj}=0\}}.$
That is, if $X$ and $Y$ are both NSND matrices, then node $i$ prefers $X$ to $Y$ if, under $X$, it has more precarious enemies.
Definition 4 (Preferable adjustment)
Given an NSND strategy matrix $X\in\Omega$ and country $i$, the preferable adjustment of $X$ with respect to $i$, denoted by $\mathrm{PA}^{i}(X)=\big(\mathrm{PA}^{i}_{kl}(X)\big)_{n\times n}$, is a strategy matrix constructed as follows:
  1. 1.
    If $p_{i}\geq\sum_{j\in\mathcal{A}_{i}}(x_{jj}+x_{ji})$, let $\mathrm{PA}^{i}(X)$ be equal to $X$ except for the changes to the following entries:
    1. (a)
      $\mathrm{PA}^{i}_{ii}(X)=p_{i}-\sum_{j\in\mathcal{A}_{i}}(x_{jj}+x_{ji})$;
    2. (b)
      $\mathrm{PA}^{i}_{ij}(X)=\mathrm{PA}^{i}_{ji}(X)=x_{jj}+x_{ji}$ and $\mathrm{PA}^{i}_{jj}(X)=0$, for any $j\in\mathcal{A}_{i}$.
  2. 2.
    If $p_{i}<\min_{j\in\mathcal{A}_{i}}(x_{jj}+x_{ji})$, let $\mathrm{PA}^{i}(X)$ be equal to $X$ except for the changes to the following entries:
    1. (a)
      Randomly pick a node $r$ from $argmin_{j\in\mathcal{A}_{i}}(x_{jj}+x_{ji})$. Let $\mathrm{PA}^{i}_{ii}(X)=0$, $\mathrm{PA}^{i}_{ir}(X)=\mathrm{PA}^{i}_{ri}(X)=p_{i}$, and $\mathrm{PA}^{i}_{rr}(X)=x_{rr}+x_{ri}-p_{i}$;
    2. (b)
      For any $j\in\mathcal{A}_{i}\setminus\{r\}$, $\mathrm{PA}^{i}_{ij}(X)=\mathrm{PA}^{i}_{ji}(X)=0$ and $\mathrm{PA}^{i}(X)_{jj}=x_{jj}+x_{ji}$.
  3. 3.
    If $\min_{j\in\mathcal{A}_{i}}(x_{jj}+x_{ji})\leq p_{i}<\sum_{j\in\mathcal{A}_{i}}(x_{jj}+x_{ji})$, let $k=\text{card}(\mathcal{A}_{i})$ and index the nodes in $\mathcal{A}_{i}$ as $j_{1},\dots,j_{k}$, with $x_{j_{1}j_{1}}+x_{j_{1}i}\leq x_{j_{2}j_{2}}+x_{j_{2}i}\leq\dots\leq x_{j_{k}j_{k}}+x_{j_{k}i}$. In this scenario, there exists $m\in\{1,\dots,k-1\}$ such that
    $\sum_{s=1}^{m}(x_{j_{s}j_{s}}+x_{j_{s}i})\leq p_{i}<\sum_{s=1}^{m+1}(x_{j_{s}j_{s}}+x_{j_{s}i}).$
    Let $\mathrm{PA}^{i}(X)$ be equal to $X$ except for the changes to the following entries: $\mathrm{PA}^{i}_{ii}(X)=0$ and,
    1. (a)
      for any $s\in\{1,\dots,m\}$, $\mathrm{PA}^{i}_{ij_{s}}(X)=\mathrm{PA}^{i}_{j_{s}i}(X)=x_{j_{s}j_{s}}+x_{j_{s}i}$ and $\mathrm{PA}^{i}_{j_{s}j_{s}}(X)=0$;
    2. (b)
      $\mathrm{PA}^{i}_{ij_{m+1}}(X)=\mathrm{PA}^{i}_{j_{m+1}i}(X)=p_{i}-\sum_{s=1}^{m}(x_{j_{s}j_{s}}+x_{j_{s}i})$ and $\mathrm{PA}^{i}_{j_{m+1}j_{m+1}}(X)=x_{j_{m+1}j_{m+1}}+x_{j_{m+1}i}-\mathrm{PA}^{i}_{j_{m+1}i}(X)$;
    3. (c)
      for any $s\in\{m+2,\dots,k\}$, $\mathrm{PA}^{i}_{ij_{s}}(X)=\mathrm{PA}^{i}_{j_{s}i}(X)=0$ and $\mathrm{PA}^{i}_{j_{s}j_{s}}(X)=x_{j_{s}j_{s}}+x_{j_{s}i}$.
Although the construction of the preferable adjustment appears complex, its underlying intuition is straightforward: by reallocating power, country $i$ seeks to maximize the number of its precarious antagonistic countries. Meanwhile, each antagonistic country of $i$ simultaneously adjusts its power allocations on $i$ and itself to preserve the symmetry of $\mathrm{PA}^{i}(X)$. Consequently, $\mathrm{PA}^{i}(X)$ remains an NSND strategy matrix, and its $i$-th row represents country $i$’s best response to $X$. The above reasoning is formalized as follows, and we refer to an earlier conference version of this paper  [zhangNashequilibriumSeekingAlgorithm2023] for the proof.
Lemma 1 (Properties of preferable adjustment)
Given any NSND strategy matrix $X\in\Omega$ and any $i\in\mathcal{V}$,
  1. 1.
    $\mathrm{PA}^{i}(X)$ is also an NSND strategy matrix;
  2. 2.
    for any $Y\in\Omega$ identical to $X$ except possibly in the $i$-th row, country $i$ prefers $\mathrm{PA}^{i}(X)$ to $Y$, i.e., $\mathrm{PA}^{i}(X)\geq_{i}Y$.
  3. 3.
    $\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{\mathrm{PA}^{i}_{jj}(X)=0\}}\geq\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}=0\}}$.
Given all the preparation work above, we are ready to prove the existence of Nash equilibrium as stated in Theorem 1.
Proof of Theorem 1: Given $n$ countries with powers $p_{1},\dots,p_{n}$ and their respective sets of friendly and antagonistic countries $\mathcal{F}_{1},\mathcal{A}_{1},\dots,\mathcal{F}_{n},\mathcal{A}_{n}$, we construct an iteration process $\{X(t)\}_{t\in\mathbb{N}}$ with $X(0)=\text{diag}(p_{1},\dots,p_{n})$. By Definition 3, $X(0)$ is an NSND strategy matrix. The iteration of $X(t)$ is conducted via the following two procedures.
Procedure 1: Given any NSND strategy matrix $X(t)$, randomly pick a pair $(i,j)$ satisfying: 1) $j\in\mathcal{A}_{i}$; 2) $x_{ii}(t)\neq 0$ and $x_{jj}(t)\neq 0$. Then let $X(t+1)=\mathrm{PA}^{i}(X(t))$. Repeat the above process until there does not exist such pair $(i,j)$.
Procedure 2: Given any NSND strategy matrix $X(t)$, let $\mathcal{B}=\{i\in\mathcal{V}\,|\,x_{ii}(t)=0\}$. Randomly pick an $i\in\mathcal{B}$ such that $\mathrm{PA}^{i}(X(t))$ satisfies
$\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{\mathrm{PA}^{i}_{jj}(X(t))=0\}}\geq\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t)=0\}}+1.$
Then let $X(t+1)=\mathrm{PA}^{i}(X(t))$ and return to Procedure 1. If no such $i\in\mathcal{B}$ exists, then terminate the iteration process.
Since $X(0)$ is an NSND strategy matrix and all the iterations in Procedure 1 and 2 are preferable adjustments, $X(t)$ is NSND for any $t\geq 0$. For any $X\in\Omega$, define
$V(X)=\sum_{i=1}^{n}\mathbf{1}_{\{x_{ii}=0\}}.$
At any $t$, suppose $X(t+1)$ is obtained from $X(t)$ via a preferable adjustment in Procedure 1 for some $i$, then, according to the conditions for triggering Procedure 1, $x_{ii}(t)\neq 0$ and there exist $j\in\mathcal{A}_{i}$ such that $x_{jj}(t)\neq 0$. In this case,
  1. 1.
    if $p_{i}\geq\sum_{j\in\mathcal{A}_{i}}(x_{jj}(t)+x_{ji}(t))$, then, by Definition 4, we have $x_{ii}(t+1)\geq 0$ and, for any $j\in\mathcal{A}_{i}$,
    $x_{ij}(t+1)$ $=x_{ji}(t+1)=x_{jj}(t)+x_{ji}(t),$
    $x_{jj}(t+1)$ $=0.$
    In addition, for any $k\!\notin\!\mathcal{A}_{i}\!\cup\!\{i\}$, $x_{kk}(t\!+\!1)\!=\!x_{kk}(t)$. Since $x_{ii}(t)\neq 0$, $x_{ii}(t+1)\geq 0$, $x_{jj}(t+1)=0$ for any $j\in\mathcal{A}_{i}$, and there exists $j\in\mathcal{A}_{i}$ such that $x_{jj}(t)\neq 0$, we have
    $V$ $(X(t+1))-V(X(t))$
    $=\sum_{j\in\mathcal{A}_{i}\cup\{i\}}\mathbf{1}_{\{x_{jj}(t+1)=0\}}-\sum_{j\in\mathcal{A}_{i}\cup\{i\}}\mathbf{1}_{\{x_{jj}(t)=0\}}\geq 1.$
  2. 2.
    if $p_{i}<\sum_{j\in\mathcal{A}_{i}}(x_{jj}(t)+x_{ji}(t))$, then, by Definition 4,
    1. (a)
      $x_{ii}(t+1)=0$ (while $x_{ii}(t)\neq 0$);
    2. (b)
      $x_{kk}(t+1)=x_{kk}(t)$ for any $k\notin\mathcal{A}_{i}\cup\{i\}$.
    In addition, according to Lemma 1, we have $X(t+1)\geq_{i}X(t)$, which, according to Fact 2, implies that $\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t+1)=0\}}\geq\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t)=0\}}$. Therefore, we have
    $V$ $(X(t+1))-V(x(t))$
    $=\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t+1)=0\}}-\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t+1)=0\}}$
    $\quad+\mathbf{1}_{\{x_{ii}(t+1)=0\}}-\mathbf{1}_{\{x_{ii}(t)=0\}}\geq 1.$
Combining 1) and 2), we conclude that, after each iteration via Procedure 1, $V(X(t+1))=V(X(t))+1$.
Now we consider the case when $X(t+1)$ is obtained from $X(t)$ via an preferable adjustment in Procedure 2 for some $i$. Since Procedure 2 requires $x_{ii}(t)=0$ and since $X(t)$ is an NSND strategy matrix, we have $p_{i}=\sum_{j\in\mathcal{A}_{i}}x_{ij}(t)=\sum_{j\in\mathcal{A}_{i}}x_{ji}(t)$. Suppose $x_{jj}(t)=0$ for any $j\in\mathcal{A}_{i}$, by Definition 4,
$x_{ij}(t+1)=x_{ij}(t)=x_{ji}(t)=x_{ji}(t+1)$
and $x_{ii}(t+1)=0$. In this case,
$\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t+1)=0\}}=\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t)=0\}},$
which contradicts the rule of Procedure 2. Therefore, there must exist $j\in\mathcal{A}_{i}$ such that $x_{jj}(t)>0$, which leads to
$p_{i}<\sum_{j\in\mathcal{A}_{i}}(x_{jj}(t)+x_{ji}(t)).$
By Definition 4, $x_{ii}(t+1)=0$ and $x_{kk}(t+1)=x_{kk}(t)$ for any $k\notin\mathcal{A}_{i}\cup\{i\}$. Therefore,
$V($ $X(t+1))-V(x(t))$
$=\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t+1)=0\}}-\sum_{j\in\mathcal{A}_{i}}\mathbf{1}_{\{x_{jj}(t)=0\}}\geq 1.$
Based on the above discussions for both Procedure 1 and Procedure 2, we conclude that after each iteration of $X(t)$, $V(X(t))$ increases by at least 1. Moreover, since $V(X(t))$ is upper bound by $n$, the iteration $X(t)$ must terminate at some finite time step $T$.
Now we proceed to prove that $X(T)$ must be a Nash equilibrium of the static power-allocation game. For any $i$, suppose $x_{ii}(T)\neq 0$. Since Procedure 1 can no longer be activated, we have $x_{jj}(T)=0$ for any $j\in\mathcal{A}_{i}$. According to Fact 1, all of $i$’s antagonistic countries are already precarious. That is, $i$ cannot further increase the number of its precarious antagonistic countries. Moreover, since $X(T)$ is NSND, $i$ and all of $i$’s friendly countries are either safe or precarious under $X(T)$. Therefore, the $i$-th row of $X(T)$ is already country $i$’s best response to $X(T)$. Suppose $x_{ii}(T)=0$. Since Procedure 2 cannot be activated, the preferable adjustment of $X(T)$ for $i$ does not increase the number of $i$’s precarious antagonistic countries. Moreover, since $X(T)$ is NSND, $i$ and all of $i$’s friendly countries are either safe or precarious under $X(T)$. Therefore, the $i$-th row of $X(T)$ is already country $i$’s best response to $X(T)$.
Now we have proved that each row of $X(T)$ is the corresponding country’s best response to $X(T)$. That is, $X(T)$ is a Nash equilibrium. Finally, since $X(T)$ is an NSND strategy matrix, no country under $X(T)$ is unsafe and no country supports any of its friendly countries, i.e., $x_{ij}(T)=0$ for any i,ji,\,j such that $j\in\mathcal{F}_{i}\setminus\{i\}$. This concludes the proof.
Theorem 1 shows that, for any static power-allocation game defined in Section 2.1, there exists a special Nash equilibrium in which every country remains non-dangerous without providing support to others. Interpreting positive power allocation between two countries as an alliance relationship, this equilibrium conceptually resonates with, and in some sense rationalizes, the real-world political agenda known as the “Non-Aligned Movement” [mitalNonalignedMovementIts].
2.3 Extreme-case analysis: Nash equilibria in a fully antagonistic world
In this subsection, we consider an extreme case where all countries are mutually antagonistic. The following theorem determines the maximum number of countries that can remain safe in such a world.
Theorem 2 (Number of safe countries in a fully antagonistic world)
Consider the static power-allocation game defined in Section 2.1 with $n\geq 3$ countries. Suppose that each country has a positive power and all the countries are hostile to each other, i.e., $\mathcal{A}_{i}=\mathcal{V}\setminus\{i\}$ for any $i\in\mathcal{V}$.
  1. 1.
    Given any power vector $p\in\mathbb{N}^{n}$ and at any Nash Equilibrium $X$, there are at most n1n-1 safe countries. Namely, it is impossible for every country to be safe.
  2. 2.
    If there exists a country $i\in\mathcal{V}$ such that $p_{i}>\sum_{j\neq i}p_{j}$, then there is only one safe country at any Nash Equilibrium, which is country $i$.
  3. 3.
    If $p_{i}\leq\sum_{j\neq i}p_{j}$ for all $i\in\mathcal{V}$, then there always exists a Nash Equilibrium with no safe country.
Proof 2.3.
We first prove Statement 1) by contradiction. Suppose that all countries are safe at some Nash Equilibrium $X\in\Omega$, which implies that no country could be in a more preferred situation by reallocating its power. Then, for any country $i,j\in\mathcal{V}$ with $i\neq j$, country $j$ cannot make country $i$ unsafe by assigning all $j$’s power on attacking $i$. That is,
$p_{j}\!-\!x_{ji}<\!\sum_{s\in\mathcal{F}_{i}}x_{si}+\!\!\sum_{s\in\mathcal{A}_{i}}x_{is}\!-\!\!\sum_{s\in\mathcal{A}_{i}}x_{si}=x_{ii}\!+\!\sum_{s\neq i}(x_{is}\!-\!x_{si}).$
Let $j=i+1$. Here we treat node $n+1$ as node $1$ and treat node 1-1 as node $n$. By summing the left-hand side of the above inequality over all $i\in\mathcal{V}$, we obtain
$\sum_{i\in\mathcal{V}}p_{i}-\!\sum_{i\in\mathcal{V}}x_{ii-1}\!<\!\sum_{i\in\mathcal{V}}x_{ii}\!+\!\sum_{i\in\mathcal{V}}\sum_{s\neq i}(x_{is}\!-\!x_{si})=\!\sum_{i\in\mathcal{V}}x_{ii}.$
Hence,
$\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}x_{ij}=\sum_{i\in\mathcal{V}}p_{i}<\sum_{i\in\mathcal{V}}(x_{ii-1}+x_{ii}),$
which leads to a contradiction and thus proves Statement 1).
The proof of Statement 2) is straightforward. If there exists $i\in\mathcal{V}$ with $p_{i}>\sum_{j\neq i}p_{j}$, no matter how other countries allocate their powers, it is impossible to make country $i$ unsafe. In the meanwhile, given the power-allocation strategies of other countries, country $i$ can make all countries except itself unsafe by assigning power $p_{j}$ on attacking country $j$ for any $j\in\mathcal{V}$. Therefore, at any Nash equilibrium, there is only one safe country, which is country $i$.
Now we prove Statement 3). We first show by induction that, when $n\geq 3$ and $\sum_{i\in\mathcal{V}}p_{i}$ is an even integer, there exists a Nash equilibrium such that every country is precarious and thus no one is safe. When $n=3$, we construct a strategy matrix $X^{*}=(x^{*}_{ij})_{3\times 3}$ as follows:
$x^{*}_{11}$ $\!=\!x^{*}_{22}\!=\!x^{*}_{33}\!=\!0;\,\,$ $x^{*}_{12}\!=\!x^{*}_{21}\!=\!\frac{p_{1}+p_{2}-p_{3}}{2};$ (1)
$x^{*}_{13}$ $\!=\!x^{*}_{31}\!=\!\frac{p_{1}+p_{3}-p_{2}}{2};$ $x^{*}_{23}\!=\!x^{*}_{32}\!=\!\frac{p_{2}+p_{3}-p_{1}}{2}.$
Since $p_{i}\leq\sum_{j\neq i}p_{j}$ for any $i\in\mathcal{V}$, $X^{*}$ is entry-wise non-negative. As $p_{1}+p_{2}+p_{3}$ is even, any off-diagonal entry of $X^{*}$ is an integer. Moreover, $\sum_{j\in\mathcal{V}}x^{*}_{ij}=p_{i}$ for any $i\in\mathcal{V}$, so $X^{*}$ is a feasible strategy matrix. Moreover, one can check that every country is precarious in this case, which is already the most preferred situation for every country. Therefore, such a strategy matrix $X^{*}$ is a Nash equilibrium.
Suppose that the induction hypothesis holds for some $n\geq 3$. In the case with $n+1$ countries, without loss of generality, let $p_{1}\leq p_{2}\leq...\leq p_{n+1}$. We have $p_{n+1}\leq p_{1}+p_{2}+...+p_{n}$. Our discussion is split into two cases:
Case 1: $p_{n+1}+2p_{n-1}>\sum_{i=1}^{n}p_{i}$. We can manually construct a Nash-equilibrium strategy matrix $X^{*}=(x^{*}_{ij})_{(n+1)\times(n+1)}$ such that no country is safe. This matrix $X^{*}$ is constructed as follows:
$x^{*}_{i,n+1}$ $=x^{*}_{n+1,i}=p_{i},\quad\text{for any }i<n-1;$
$x^{*}_{n-1,n}$ $=x^{*}_{n,n-1}=\frac{\sum_{i=1}^{n}p_{i}-p_{n+1}}{2};$
$x^{*}_{n,n+1}$ $=x^{*}_{n+1,n}\!=\!\frac{p_{n+1}\!+\!2p_{n-1}\!-\!\!\sum_{i=1}^{n}p_{i}}{2}\!+\!p_{n}\!-\!p_{n-1};$
$x^{*}_{ij}$ $=x^{*}_{ji}=0,\quad\text{otherwise}.$
Since $p_{n+1}\leq\sum_{i=1}^{n}p_{i}$, each entry of $X^{*}$ is non-negative. Since $2x^{*}_{n,n-1}+2p_{n+1}=\sum_{i=1}^{n+1}p_{i}$ is even, $x^{*}_{n,n-1}=x^{*}_{n-1,n}$ is an integer. Since $2x^{*}_{n,n+1}+\sum_{i=1}^{n+1}p_{i}=2(p_{n-1}+p_{n+1})$ is even, $x^{*}_{n,n+1}=x^{*}_{n+1,n}$ is an integer. As a result, each entry of $X^{*}$ is an integer. In addition, via simple calculations, one can check that $\sum_{j=1}^{n+1}x^{*}_{ij}=p_{i}$ for any $i\in\{1,\dots,n,n+1\}$. That is, $X^{*}$ is a feasible strategy matrix. Moreover, since all the countries are antagonistic to each other and the diagonals of $X^{*}$ are all zero, the symmetry of $X^{*}$ determines that every country is precarious under $X^{*}$, which in turn implies that $X^{*}$ is a Nash equilibrium, where no country is safe.
Case 2: $p_{n+1}-(p_{n}-p_{n-1})\leq\sum_{i=1}^{n-2}p_{i}$. Then there exists $1\leq k\leq n-2$ such that
$\sum_{i=1}^{k-1}p_{i}<p_{n+1}-(p_{n}-p_{n-1})\leq\sum_{i=1}^{k}p_{i}.$
Here we adopt the shorthand $\sum_{i=1}^{0}p_{i}=0$ if needed. We construct an auxiliary system with $n$ mutually antagonistic countries whose powers $p_{1}^{*},\dots,p_{n}^{*}$ are given as follows:
$p_{i}^{*}$ $=0,\quad\text{for any }i<k;$
$p_{k}^{*}$ $=\sum_{j=1}^{k}p_{j}-p_{n-1}+p_{n}-p_{n+1};$
$p_{i}^{*}$ $=p_{i},\quad\text{for any }k<i\leq n-1;$
$p_{n}^{*}$ $=p_{n-1}.$
Since $p_{n+1}-(p_{n}-p_{n-1})\leq\sum_{i=1}^{n-2}p_{i}$, for any $i\in\{1,\dots,n\}$, $p_{i}^{*}$ is a non-negative integer. In addition,
$\sum_{i=1}p^{*}_{i}=\sum_{j=1}^{n-1}p_{j}+p_{n}-p_{n+1}=\sum_{i=1}^{n+1}p_{i}-2p_{n+1}.$
Therefore, $\sum_{i=1}^{n}p_{i}^{*}$ is even. The order $p_{1}\leq p_{2}\leq\dots\leq p_{n}$ implies $p^{*}_{1}\leq p^{*}_{2}\leq\dots\leq p^{*}_{n}$. Furthermore,
$2p^{*}_{n}-\sum_{j=1}^{n}p_{j}^{*}=p_{n+1}-(p_{n}-p_{n-1})-\sum_{j=1}^{n-2}p_{i}\leq 0,$
and hence $2p^{*}_{i}\leq\sum{j=1}^{n}p_{j}^{*}$ for all $i\in\{1,\dots,n\}$. Thus, all conditions for the induction hypothesis are satisfied. Consequently, for this auxiliary system, there exists a Nash equilibrium $X^{*}=(x^{*}_{ij}){n\times n}$ under which every country is precarious.
For the original $(n+1)$-country system with powers $p_{1},\dots,p_{n+1}$, construct a strategy matrix $X=(x_{ij})_{(n+1)\times(n+1)}$ whose first $n$ rows and columns coincide with $X^{*}$. Let
$x_{n+1,i}=x_{i,n+1}=p_{i}-p_{i}^{*},\quad\text{for any }i\in\{1,\dots,n\},$
and $x_{n+1,n+1}=0$. One can verify that $X$ is feasible, and that every country is precarious under $X$. Hence, $X$ is a Nash equilibrium. This completes the induction proof that, when the sum of all countries’ powers is even, there exists a Nash equilibrium such that all countries are precarious.
If the sum of all countries’ power is odd, take any three countries with positive power, denoted by $i_{1},i_{2},i_{3}$, and temporarily reduce their power by 1. Now the countries’ total power is even. We have proved that, in this case, there exists a Nash Equilibrium $X$ such that all countries are precarious. Then we return the reduced powers to the three countries and add $1$ to $x_{i_{1}i_{2}},x_{i_{2}i_{3}},and\ x_{i_{3}i_{1}}$ respectively. The new strategy matrix remains feasible. Moreover, it is still a Nash equilibrium since all countries are precarious. This completes the proof of Statement 3).
2.4 Inverse Inference of Unknown Edge Weights in Network Connections
In complex network systems, the strength of connections between nodes is often heterogeneous. Varying edge weights reflect the diversity of interactions among nodes. Such disparities in weighting not only influence the evolutionary dynamics of the system as a whole, but also largely determine the role and functional importance of each node within the network.
In this context, inverse inference plays an indispensable role. Since these weights often cannot be directly observed or measured, relying solely on prior information is insufficient for their accurate characterization. Therefore, it is necessary to perform inverse inference based on observed data. Through inverse inference, one can recover the hidden distribution of weights from Nash equilibrium trajectories, thereby extracting the actual intensity of interactions between nodes. More importantly, inverse inference not only reveals the true relationships among nodes but also enhances the predictive power and interpretability of the model, ensuring that research outcomes remain robust in more complex and uncertain environments.
Building upon the static power-allocation game model, we further extend our study to a more general form of influence weights between nodes. We modify the preference axiom (2) to the following form:
$\sum_{j\in E_{i}}w_{ij}I_{s_{j}(X)\leq 0}$ $+\sum_{l\in A_{i}}w_{il}I_{s_{l}(X)\geq 0}\geq$
$\sum_{j\in E_{i}}w_{ij}I_{s_{j}(Y)\leq 0}+\sum_{l\in A_{i}}w_{il}I_{s_{l}(Y)\geq 0},$
where the elements $w_{ij}$ of theweight matrix $\mathbf{W}$ denotes the importance of node $j$ to node $i$. Our objective is to utilize a series of observed approximate Nash equilibrium strategy matrices from real-world data to inversely infer the differences in weights between nodes. In the following, we propose an optimization algorithm to address this problem through detailed analysis and the inspiration mainly comes from [paganGameTheoreticalInference2019a].
Suppose we have an observed matrix $\mathbf{X}^{*}$. For any node $i$, we aim to deduce the influence weights $\mathbf{w}_{i}=(w_{i1},w_{i2},\dots,w_{in})$ from other nodes on node $i$. Let $\mathbf{X}=(\mathbf{x}_{i},\mathbf{x}_{-i})$, where $\mathbf{x}_{i}=(x_{i1},x_{i2},\dots,x_{in})$ represents the strategy chosen by node $i$, and $\mathbf{x}_{-i}$ denotes the strategy profile of all other nodes. Define $\mathcal{X}_{i}=\{\mathbf{x}_{i}\in\mathbb{R}^{m}\mid\sum_{j=1}^{m}x_{ij}=p_{i}\}$ as the set of all feasible strategies for node $i$. The utility of node $i$ is given by:
$V_{i}(\mathbf{x}_{i},\mathbf{x}_{-i},\mathbf{w}_{i})=\sum_{j=1}^{m}w_{ij}I_{s_{i}(\mathbf{X})\geq 0}$ $+\sum_{j\in A_{i}}w_{ij}I_{s_{j}(\mathbf{X})\geq 0}$
$+\sum_{j\in E_{i}}w_{ij}I_{s_{j}(\mathbf{X})\leq 0}.$
According to variational inequality theory, when the strategy profile of all nodes constitutes a Nash equilibrium, no node $i$ can unilaterally change its own strategy to improve its utility. For a Nash equilibrium $\mathbf{X}^{*}=(\mathbf{x}_{i}^{*},\mathbf{x}_{-i}^{*})$, define:
$e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})=V_{i}(\mathbf{x}_{i},\mathbf{x}_{-i}^{*},\mathbf{w}_{i})-V_{i}(\mathbf{x}_{i}^{*},\mathbf{x}_{-i}^{*},\mathbf{w}_{i}).$
Note that $e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ is linear with respect to each component of $\mathbf{w}_{i}$. For any feasible $\mathbf{x}_{i}$, $e_{i}$ should be non-positive. However, whenever the weight parameter $\mathbf{w}_{i}$ leads to a strategy $\mathbf{x}_{i}$ that violates the Nash equilibrium condition, the function $e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ takes a positive value. This positive violation can be interpreted as an error, potentially arising from factors such as the bounded rationality of nodes. Therefore, we introduce an error term defined as $e_{i}^{+}(\mathbf{x}_{i},\mathbf{w}_{i})=\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}$, representing the degree of violation of strategy $\mathbf{x}_{i}$ compared to the Nash equilibrium strategy $\mathbf{x}_{i}^{*}$ under parameter $\mathbf{w}_{i}$. A common approach to measure the distance from Nash equilibrium is to integrate the squared error term $e_{i}^{+}(\mathbf{x}_{i},\mathbf{w}_{i})$ over the entire action space, which aligns with the concept of the $L_{2}$ norm [paganGameTheoreticalInference2019a]:
$d_{i}(\mathbf{w}_{i}):=\left(\int_{\mathcal{X}_{i}}(e_{i}^{+}(\mathbf{x}_{i},\mathbf{w}_{i}))^{2}d\mathbf{x}_{i}\right)^{1/2}.$
The following lemma establishes the continuous differentiability and convexity of $d_{i}(\mathbf{w}_{i})$, providing the theoretical foundation for finding its minimum.
Lemma 2.4.
(Continuous Differentiability and Convexity of $d_{i}(\mathbf{w}_{i})$) When $\mathbf{w}_{i}$ takes values in a bounded set $\mathcal{W}$, $d_{i}(\mathbf{w}_{i}):\mathcal{W}\subseteq\mathbb{R}^{m}\to\mathbb{R}$ is a continuously differentiable convex function.
Proof 2.5.
First, assume that all components in $\mathcal{W}$ do not exceed $\hat{W}$. Note that for any $\mathbf{x}_{i}\in\mathbb{R}^{n}$, $\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}$ has a continuous first-order derivative with respect to $\mathbf{w}_{i}$, and its gradient is:
$\nabla_{\mathbf{w}_{i}}$ $(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}$
$=\begin{cases}0,&\text{if }e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})<0,\\ 2e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i}),&\text{otherwise}.\end{cases}$
This holds because $e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ is linear in the components of $\mathbf{w}_{i}$, making $\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ continuous with respect to $\mathbf{w}_{i}$. Define $\mathcal{X}_{+}(\mathbf{w}_{i})=\{\mathbf{x}_{i}\in\mathcal{X}\mid e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})>0\}$ and $\mathcal{X}_{-}(\mathbf{w}_{i})=\{\mathbf{x}_{i}\in\mathcal{X}\mid e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\leq 0\}$. For any $j\in\{1,2,\dots,m\}$, since $(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}$ is integrable with respect to $\mathbf{x}_{i}$ for any fixed $w_{ij}$, and
$\left|\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}\right|\leq 4m\hat{W}^{2},$
it follows from Theorem 2.27 in [follandRealAnalysis2013] that:
$\frac{\partial d_{i}(\mathbf{w}_{i})}{\partial w_{ij}}$ $=\int_{\mathcal{X}}\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}d\mathbf{x}_{i}$
$=\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}d\mathbf{x}_{i}$
$\ \ \ \ +\int_{\mathcal{X}_{-}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}d\mathbf{x}_{i}$
$=\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ij}}(e_{i}(\mathbf{x}_{i},\mathbf{w}_{i}))^{2}d\mathbf{x}_{i}$
$=\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}2e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}$
$=\int_{\mathcal{X}}2\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}.$
Note that since $e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ is linear in $\mathbf{w}_{i}$, $\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$ is independent of $\mathbf{w}_{i}$, implying that $\frac{\partial d_{i}(\mathbf{w}_{i})}{\partial w_{ij}}$ is continuous in $\mathbf{w}_{i}$. Therefore, the gradient in vector form is:
$\nabla_{\mathbf{w}_{i}}d_{i}(\mathbf{w}_{i})=\int_{\mathcal{X}}2\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}.$
Next, we compute the Hessian matrix of $d_{i}(\mathbf{w}_{i})$. Note that $\frac{\partial}{\partial w_{ij}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}$ is integrable with respect to $\mathbf{x}_{i}$ for any fixed $w_{ij}$, and for any $k\in\{1,2,\dots,m\}$, $\frac{\partial^{2}}{\partial w_{ij}\partial w_{ik}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}$ exists, with
$\left|\frac{\partial^{2}}{\partial w_{ij}\partial w_{ik}}(\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\})^{2}\right|\leq 4.$
Thus, by Theorem 2.27 in [follandRealAnalysis2013]:
$\frac{\partial^{2}d_{i}(\mathbf{w}_{i})}{\partial w_{ij}\partial w_{ik}}$
$\ \ \ \ \ =\frac{\partial}{\partial w_{ik}}\int_{\mathcal{X}}2\max\{0,e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\}\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}$
$\ \ \ \ \ =\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ik}}\left[2e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\right]d\mathbf{x}_{i}$
$\ \ \ \ \ =2\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}[\frac{\partial}{\partial w_{ik}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})$
$\ \ \ \ \ \ \ \ \ +e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial^{2}}{\partial w_{ij}\partial w_{ik}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})]d\mathbf{x}_{i}$
$\ \ \ \ \ =2\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\frac{\partial}{\partial w_{ik}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\frac{\partial}{\partial w_{ij}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})d\mathbf{x}_{i}.$
Therefore, the Hessian matrix is $H(d_{i}(\mathbf{w}_{i}))=2\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})(\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i}))^{T}d\mathbf{x}_{i}$, which is positive semi-definite because for any $\mathbf{y}\in\mathbb{R}^{n}$:
$\mathbf{y}^{T}H(d_{i}(\mathbf{w}_{i}))\mathbf{y}=2\int_{\mathcal{X}_{+}(\mathbf{w}_{i})}\left(\mathbf{y}^{T}\nabla_{\mathbf{w}_{i}}e_{i}(\mathbf{x}_{i},\mathbf{w}_{i})\right)^{2}d\mathbf{x}_{i}\geq 0.$
Hence, $d_{i}(\mathbf{w}_{i})$ is a convex function.
From Lemma 3, $d_{i}(\mathbf{w}_{i})$ is a continuously differentiable convex function on any bounded set. This allows us to propose the following meaningful optimization problem:
Problem 2.6.
(Minimum Nash Equilibrium Distance Problem) Given a signed network with $n$ nodes, for all nodes $i$, find the weight parameter $\hat{\mathbf{w}}_{i}$ that solves the following optimization problem:
$\min_{\mathbf{w}_{i}\in\mathcal{W}}(d(\mathbf{w}_{i}))^{2}$
$\text{s.t. }\mathbf{w}_{i}\in\mathcal{W}.$
Although the objective function is convex, the lack of guaranteed strict convexity makes direct optimization challenging. We first approximate the Riemann integral by discretizing the feasible strategy space into a set of discrete points. Suppose we select feasible strategies $\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2},\dots,\mathbf{x}_{i}^{n_{i}}$. The optimization problem then transforms into:
Problem 2.7.
(Discrete Minimum Nash Equilibrium Distance Problem) Given a signed network with $n$ nodes, for all nodes $i$, find the weight parameter $\hat{\mathbf{w}}_{i}$ that solves the following optimization problem:
$\min_{\mathbf{w}_{i}\in\mathcal{W}}\sum_{j=1}^{n_{i}}(e_{i}^{(j,+)}(\mathbf{w}_{i}))^{2}$
$\text{s.t. }\mathbf{w}_{i}\in\mathcal{W}.$
In complex network modeling and inference, if an edge weight is inferred to be non-zero, it implies a substantive interaction between the corresponding nodes. To avoid spurious connections caused by observational noise or numerical approximation, it is necessary to set a fixed positive threshold, ensuring that all non-zero weights strictly exceed this value. This constraint not only theoretically guarantees the interpretability of the edge’s existence but also effectively distinguishes genuine interactions from artifacts induced by random perturbations, thereby enhancing the scientific validity and robustness of the model results. Furthermore, the model may face overfitting issues, causing it to rely excessively on limited observed data and lose its ability to capture underlying general patterns. To address this, we introduce a regularization term in the inference process to balance model fit and complexity within the optimization objective. Consequently, we ultimately solve the following optimization problem to infer the weights influencing node $i$ via inverse inference:
Problem 2.8.
(Discrete Minimum Nash Equilibrium Distance Problem) Given a signed network with $n$ nodes, for all nodes $i$, find the weight parameter $\hat{\mathbf{w}}_{i}$ that solves the following optimization problem:
$\min_{\mathbf{w}_{i}\in\mathcal{W}}\sum_{j=1}^{n_{i}}(e_{i}^{(j,+)}(\mathbf{w}_{i}))^{2}+\|\mathbf{w}_{i}\|_{2}^{2}$
$\text{s.t. }w_{ij}\geq 0.01,\quad\forall j\in\mathcal{A}_{i}\cup\mathcal{F}_{i}\backslash\{i\}.$
Note that the first part of the objective function is linear in $\mathbf{w}_{i}$, while the second part is a convex quadratic function. As the components of $\mathbf{w}_{i}$ tend to positive infinity, the objective function also tends to positive infinity, ensuring that if a minimum exists, it must lie within a bounded set. Moreover, since the objective function is strictly convex over this bounded set and the variables lie in a convex set, the solution to this optimization problem exists and is unique.
3 Dynamic model: co-evolution of countries’ powers and strategies
Intuitively, a country in a safe environment is more likely to prosper, whereas the power of a country in danger tends to decline. This pattern is supported by empirical evidence provided in Section 4.2, which relates countries’ Nash-equilibrium states to their subsequent GDP growth. These observations motivate us to study how signed network structures shape the evolution of national powers and, when conflicts are inevitable, which network configurations better foster global prosperity. To address these questions, we extend the static power-allocation game to a dynamic setting incorporating the co-evolution of countries’ powers and strategy adjustments.
3.1 Model setup
For the dynamics model, regarding the network, the countries’ powers, and power-allocation strategies, we adopt similar model setups to the static power-allocation game. Consider $n$ countries embedded in a signed network, where $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ denote the set of friendly and antagonistic neighbors for any country $i\in\mathcal{V}$, respectively. The dynamic system starts with an initial vector of powers $p(0)=\big(p_{1}(0),\dots,p_{n}(0)\big)^{\top}\in\mathbb{N}^{n}$ and an initial power-allocation strategy matrix $X(0)=(x_{ij}(0))_{n\times n}$ with $\sum_{j\in V}x_{ij}(0)=p_{i}(0)$ for any $i\in\mathcal{V}$.
To provide a more precise characterization of the dynamic model, in this section, we adopt a specific utility function satisfying the preference axioms presented in Section 2.1. For any country $i\in\mathcal{V}$, define the utility of $i$ under a strategy matrix $X$ as
$u_{i}(X)=\sum_{j=1}^{n}u_{ji}(X)+n\mathbf{1}_{\{s_{i}(X)\geq 0\}},$
where
$u_{ji}(X)=\begin{cases}1,&\text{if }j\in F_{i}\text{ and }s_{j}(X)\geq 0,\\ &\text{or if }j\in A_{i}\text{ and }s_{j}(X)\leq 0,\\ 0,&\text{otherwise}.\end{cases}$
and define
$u_{i}(X)=\sum_{k=1}^{n}u_{ki}(X)+n\mathbf{1}_{\{s_{i}(X)\geq 0\}},$
where $i,j\in\mathcal{V}$ and $X\in\mathbf{S}$.
Namely, “being not in danger” is the prioritized goal of any country. Besides that, a country seeks to maximize the number of friendly countries that are not in danger, plus the number of antagonistic countries that are not safe.
In the dynamics model, countries are also embedded in a signed network. Each country $i$ starts with an initial power $p_{i}(0)$ and an initial power allocation $(x_{i1}(0),x_{i2}(0),\ldots,x_{in}(0))$. At each time $t$, a country $i_{t}$ is randomly chosen from the set $\mathcal{V}$ to update its power and strategy. How the countries’ strategies and powers are updated is specified later in this subsection. Denote by $\{i_{t}\}$ the update sequence. Supporse that the maximal possible power of a country is $K\in\mathbb{N}_{+}$. Let $\mathbf{S}=\{X=(x_{ij})_{n\times n}\in\mathbb{N}^{n\times n}|\forall 1\leq i\leq n,\sum_{j=1}^{n}x_{ij}\leq K\}$. That is, any $X\notin\mathbf{S}$ is not a feasible strategy matrix. Let $\mathbf{X}(X,\{i_{t}\})=\{X_{t}\}_{t=0}^{\infty}$ represent the sequence of the strategy matrices starting from $X_{0}=X$ and along with the update sequence $\{i_{t}\}$. Let $\mathbf{X}(X,\{i_{t}\}_{1\leq t\leq T})=\{X_{t}\}_{t=0}^{T}$ represent the sequence of strategy matrcies of finite length. Given the strategy matrix $X$, let $s_{i}(X)=\sum_{j\in\mathcal{F}_{i}}x_{ji}+\sum_{j\in\mathcal{A}_{i}}x_{ij}-\sum_{j\in\mathcal{A}_{i}}x_{ji}$, and classify each country into one of three states: safe, precarious, or dangerous accoding to whether $s_{i}(X)$ is positive, zero, or negative. Let $p_{i}(t)$ be the power of country $i$ at time $t$, which is non-negative with a universal upper bound $K\in\mathbf{N}_{+}$. Rigorously speaking, variables like $p_{i}(t)$ depend on the initial condition and the update sequence. But, without causing any confusion, these arguments are omitted for simplicity of notations. For any set $\mathbf{D}$, denote by $N_{p}(\mathbf{D},X)$ the number of precarious countries in the set $\mathbf{D}$ with the strategy matrix $X$.
Now we specify how the countries’ powers and strategies evolve. Suppose the update sequence is $\{i_{t}\}$. At each time $t$, the update process consists of two stages. Firstly, the power of $i_{t}$ evolves according to the following rule, dependent on the state of $i_{t}$:
(a). If country $i_{t}$ is safe, i.e., $s_{i_{t}}(X_{t-1})>0$, then
$p_{i_{t}}(t)=\min\{K,p_{i_{t}}(t-1)+1\}.$
(b). If country $i_{t}$ is precarious, i.e., $s_{i_{t}}(X_{t-1})=0$, $p_{i_{t}}(t)=p_{i_{t}}(t-1).$
(c). If country $i_{t}$ is dangerous, i.e., $s_{X_{t-1}}(t-1)<0$,
$p_{i_{t}}(t)=\max\{0,p_{i_{t}}(t-1)-1\}.$
Secondly, given the strtegy matrix $X_{t-1}$, let the matrix obtained after country $i_{t}$ updates be selected uniformly at random from the set $H(X_{t-1},i_{t})$, which satisfies the following rules:
Rule 1. Country $i_{t}$ selects the strategy without ”wasting any power”. Let $\Omega$ be the set consisting of all matrices $X=(x_{ij})_{n\times n}$ that satisfy the following three conditions:
  1. [label=]
  2. 1.
    (a) $x_{i_{t}j}\leq x_{i_{t}j}(t-1)$, if $u_{ji_{t}}(X_{t-1})=1$,
  3. 2.
    (b) $x_{i_{t}j}\leq x_{i_{t}j}(t-1)+|s_{j}(t-1)|$, if $u_{ji_{t}}(X_{t-1})=0$.
  4. 3.
    (c) $x_{jk}=x_{jk}(t-1)$, for all $j\neq i_{t},k\in\mathcal{V}$.
Let $H(X_{t-1},i_{t})\subseteq\Omega$.
Rule 2. Country $i_{t}$ prioritizes the stratrgies that increase its utility. If $\Omega_{1}=\Big\{X\in\Omega\Big|u_{i_{t}}(X)>u_{i_{t}}(X_{t-1})\Big\}$ is not empty, let $H(X_{t-1},i_{t})=\Omega_{1}$.
Rule 3. Country $i_{t}$ prefers the strategies that reduce conflicts. Assume that $\Omega_{1}$ is empty. If $\Omega_{2}=\Big\{X\in\Omega\Big|u_{i_{t}}(X)\geq u_{i_{t}}(X_{t-1}),x_{ij}\leq x_{ij}(t-1),\forall j\in\mathcal{A}_{i},\text{ and }x_{ij}=x_{ij}(t-1),\forall j\in\mathcal{F}_{i}\setminus\{i\}\Big\}$ is not empty, let $H(X_{t-1},i_{t})=\Omega_{2}$. If $\Omega_{2}$ is empty and $\Omega_{3}=\Big\{X\in\Omega\Big|u_{i_{t}}(X)=u_{i_{t}}(X_{t-1})\Big\}$ is not empty, let $H(X_{t-1},i_{t})=\Omega_{3}$.
Rule 4. Country $i_{t}$ will randomly choose a feasible strategy if all its feasible strategies lead to a decrease of its utility. If $\Omega_{1},\Omega_{2}$ and $\Omega_{3}$ are all empty, let $H(X_{t-1},i_{t})=\Omega$.
The process continues until the strategy matrix remains constant no matter which country is activated, i.e., when the system reaches an equilibrium.
3.2 Convergence analysis
In this section, we prove that any dynamic game satisfying the definition in Section 3.1 will ultimately converge to an equilibrium, which is also a Nash equilibrium strategy matrix defined in Section 2.1. The theorem is presented formally as follows:
Theorem 3.9 (Convergence of the dynamic process).
Consider the dynamic processes of power evolution and reallocation defined in Section 3.1. Given any initial condition $p(0)\in\{1,2,...,K\}^{n}$ and $X_{0}$ compatible with $p(0)$, the countries’ power $p(t)$ and the strategy matrix $X_{t}$ almost surely converge to a steady state. Furthermore, the resulting steady-state strategy matrix constitutes a Nash Equilibrium of the power-allocation game defined in Section 2.1.
Before proving the theorem, some useful lemmas are needed.
Lemma 3.10.
Consider the dynamics defined in Section 3.1. If, starting from any strategy matrix $X\in\mathbf{S}$, there exists an update sequence $\mathbf{X}(X,\{i_{t}\})$, along which the strategy matrix reaches a steady state at time step $T_{X}$, then the power-allocation dynamics almost surely converges to a steady state in finite time, for any initial condition $X\in\mathbf{S}$.
Since $\mathbf{S}$ is a finite set and the state transition probabilities at time $t$ only depends on $X_{t}$, the evolution of $X_{t}$ and $p(t)$ constitutes a Markov chain. Therefore, the lemma about can be proved via a similar argument to the proof of lemma 7 in [meiConvergenceConsensusDissensus2024]. The detailed proof is thus omitted here.
As indicated by lemma 2, the stochastic dynamic process defined in Section 3.1 almost surely reaches an equilibrium in finite time if, for any initial condition, we can manually design an update sequence, along with the system reaches an equilibrium. Guided by this lemma, we prove Theorem 3.9 through the following three main steps.
Step 1: Starting with any initial condition, we show that the strategy matrix can reach a set of special matrices in finite iterations.
Step 2: We prove that there is no dangerous country given the strategy matrix obtained after step 1.
Step 3: We further prove that, after step 1, no country changes its state no matter which country is chosen to update after step 1. This property guarantees the convergence of the dynamic process.
In step 1, we will keep updating non-dangerous countries. The following lemma states the consequence of such operations.
Lemma 3.11.
Consider the dynamics defined in Section 3.1. Suppose the country selected for update is safe or precarious at time $t$. If $s_{i_{t}}(X_{t-1})\geq 0$, we have $N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t})-N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t-1})\geq u_{i_{t}}(X_{t})-u_{i_{t}}(X_{t-1}).$ That is, the increase of the number of precarious countries in the set of all the countries except the activated country is not less than the increase of the utility of the activated country.
Proof 3.12.
Given any $X\in\mathbf{S}$ and any update sequence $\{i_{t}\}$, without causing any confusion, we use the shorthand notation $\{X_{t}\}$ in this proof to denote $\mathbf{X}(X,\{i_{t}\})$, i.e. the sequence of the strategy matrices starting from $X_{0}=X$ and along with the sequence $\{i_{t}\}$. For any $i\in\mathcal{V}$ and $t\in\mathbb{N}$, we partition all the countries except $i$ into three subsets:
$\mathbf{C}^{-}_{i,t}(\mathbf{X})$ $=\{j\neq i|u_{ji}(X_{t-1})=1\ \text{and}\ u_{ji}(X_{t})=0\},$
$\mathbf{C}^{+}_{i,t}(\mathbf{X})$ $=\{j\neq i|u_{ji}(X_{t-1})=0\ and\ u_{ji}(X_{t})=1\},$
$\mathbf{C}^{0}_{i,t}(\mathbf{X})$ $=\{j\neq i|u_{ji}(X_{t-1})=u_{ji}(X_{t})\}.$
[…]
$N_{p}(\mathbf{C}^{-}_{i_{t},t}(\mathbf{X}),X_{t})-N_{p}(\mathbf{C}^{-}_{i_{t},t}(\mathbf{X}),X_{t-1})$
$\geq-card(\mathbf{C}^{-}_{i_{t},t}(\mathbf{X})),$
$N_{p}(\mathbf{C}^{+}_{i_{t},t}(\mathbf{X}),X_{t})-N_{p}(\mathbf{C}^{+}_{i_{t},t}(\mathbf{X}),X_{t-1})$
$\geq card(\mathbf{C}^{+}_{i_{t},t}(\mathbf{X})),$
$N_{p}(\mathbf{C}^{0}_{i_{t},t}(\mathbf{X}),X_{t})-N_{p}(\mathbf{C}^{0}_{i_{t},t}(\mathbf{X}),X_{t-1})$ $\geq 0.$
Here we have:
$N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t})-N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t-1})$
$=N_{p}(\mathbf{C}^{-}_{i_{t},t}(\mathbf{X}),X_{t})-N_{p}(\mathbf{C}^{-}_{i_{t},t}(\mathbf{X}),X_{t-1})$
$\ \ \ \ +(N_{p}(\mathbf{C}^{+}_{i_{t},t}(\mathbf{X}),X_{t})-N_{p}(\mathbf{C}^{+}_{i_{t},t}(\mathbf{X}),X_{t-1}))$
$\ \ \ \ +(N_{p}(\mathbf{C}^{0}_{i_{t},t}(\mathbf{X}),X_{t})-N_{p}(\mathbf{C}^{0}_{i_{t},t}(\mathbf{X}),X_{t-1}))$
$\geq card(\mathbf{C}^{+}_{i_{t},t}(\mathbf{X}))-card(\mathbf{C}^{-}_{i_{t},t}(\mathbf{X}))$
$=u_{i_{t}}(X_{t})-u_{i_{t}}(X_{t-1}).$
This concludes the proof.
With all the preparations above, now we are ready to prove Theorem 3.9.
Proof of theorem 3.9: According to the update rule, when a safe country is activated to update its power and strategy, the utility of the activated country will not decrease. Moreover, according to Lemma 3.11,
$N_{p}(\mathcal{V},X_{t})-N_{p}(\mathcal{V},X_{t-1})$
$=N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t})-N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t-1})$
$\ \ \ \ \ +N_{p}(\{i_{t}\},X_{t})-N_{p}(\{i_{t}\},X_{t-1})$
$\geq u_{i_{t}}(X_{t})-u_{i_{t}}(X_{t-1})\geq 0.$
Hence, the number of precarious countries will not decrease. Suppose that a precarious country is activated. If the utility of the activated country increases by at least one, according to Lemma 3.11, we have
$N_{p}(\mathcal{V},X_{t})-N_{p}(\mathcal{V},X_{t-1})$
$=N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t})-N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t-1})$
$\ \ \ \ \ +(N_{p}(\{i_{t}\},X_{t})-N_{p}(\{i_{t}\},X_{t-1}))$
$\geq u_{i_{t}}(X_{t})-u_{i_{t}}(X_{t-1})-1\geq 0.$
In this case, the number of precarious countries will not decrease. If the utility of the activated country does not change, then this country will modify its previous power-allocation strategy by only randomly retracting some power used to attack hostile countries without changing the state of hostile countries from ’unsafe’ to ’safe’. Consequently, the activated country remains precarious, implying that the number of precarious countries will not decrease.
Since the total number of countries is finite, starting from any initial condition, the above argument implies that, if we only activate safe or precarious countries, after a finite time steps, the number of precarious countries will remain constant. Define
$\mathbf{S_{1}}=\Big\{\tilde{X}\in\mathbf{S}\Big|$ $\{X_{t}\}_{t\in\mathbb{N}}=\mathbf{X}(\tilde{X},\{i_{t}\}_{t\in\mathbb{N}^{+}})\text{ satisfies that }$
$N_{p}(\mathcal{V},X_{t})\equiv N_{p}(\mathcal{V},\tilde{X})\text{ for all }t\in\mathbb{N}^{+}$
$\text{ if }s_{i_{t}}(X_{t-1})\geq 0\text{ for any }t\in\mathbb{N}^{+}.\Big\},$
According to the definition, we know that starting from any matrix in $\mathbf{S_{1}}$, and along with any update sequence such that no dangerous country is updated, the number of precarious countries remains constant. Moreover, for any matrix $X^{(0)}\in\mathbf{S}$, there exist a sequence of strategy matrices $\{X^{(0)}_{t}\}_{0\leq t\leq T_{1}}=\mathbf{X}(X^{(0)},\{i_{t}\}_{1\leq t\leq T_{1}})$ where $s_{i_{t}}(X^{(0)}_{t-1})\geq 0$ for any $t\leq T_{1}$ and $X^{(0)}_{T_{1}}\in\mathbf{S_{1}}$.
Next, we will show that once the strategy matrix $X_{T_{1}}\in\mathbf{S_{1}}$ is obtained, if we keep updating the precarious countries, the states of all the countries will no longer change after a finite times of updates. For any $X\in\mathbf{S}$, define
$Aid(X)=\sum_{j\in\mathcal{V}}\sum_{i\in A_{j}\setminus\{j\}}x_{ij}$
i.e., the total support between friendly countries excluding the self supports. Consider the following two cases that cover all the possible scenarios when the activated country $i_{t}$ is precarious:
(a). If $u_{i_{t}}(X_{t})-u_{i_{t}}(X_{t-1})=0$, country $i_{t}$ will modify its precious power-allocation strategy by only randomly retracting the power used to strike hostile countries, provided that such retraction does not affect the utilities received from those antagonistic countries. Therefore, $Aid(X_{t})=Aid(X_{t-1})$. Given that the strategy matrix $X_{t}$ still belongs to $\mathbf{S_{1}}$, i.e. the number of precarious countries does not change, no antagonistic countries will transit from danger to precariousness after the update. Consequently, no country’s state changes.
(b). If $u_{i_{t}}(X_{t})-u_{i_{t}}(X_{t-1})>0$, then
$N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t})-N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t-1})$
$\geq u_{i_{t}}(X_{t})-u_{i_{t}}(X_{t-1})\geq 1.$
Meanwhile, the strategy matrix $X\in\mathbf{S_{1}}$ implies that
$N_{p}(\mathcal{V},X_{t})-N_{p}(\mathcal{V},X_{t-1})=0,$
which means that:
$N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t})-N_{p}(\mathcal{V}\setminus\{i_{t}\},X_{t-1})$
$=u_{i_{t}}(X_{t})-u_{i_{t}}(X_{t-1})=1,$
and $s_{i_{t}}(X_{t})>0$.
Note that,
$s_{i}(X_{t})=\sum_{j\in\mathcal{F}_{i}}x_{ji}(t)+\sum_{j\in\mathcal{A}_{i}}x_{ij}(t)-\sum_{j\in\mathcal{A}_{i}}x_{ji}(t)$
$=p_{i}(t)+\sum_{j\in\mathcal{F}_{i},j\neq i}x_{ji}(t)-\sum_{j\in\mathcal{A}_{i}}x_{ji}(t)-\sum_{j\in\mathcal{F}_{i},j\neq i}x_{ij}(t).$
Therefore,
$s_{i_{t}}(X_{t})-s_{i_{t}}(X_{t-1})$ $=\sum_{\begin{subarray}{c}j\in\mathcal{F}_{i_{t}}\\ j\neq i_{t}\end{subarray}}x_{i_{t}j}(t)\ -\sum_{\begin{subarray}{c}j\in\mathcal{F}_{i_{t}}\\ j\neq i_{t}\end{subarray}}x_{i_{t}j}(t)$
$=Aid(X_{t-1})-Aid(X_{t}).$
In this scenario, the decrease of $a(t)$ is equal to the increase of $s_{i_{t}}(X_{t})$, which is a negative integer. Moreover, because for all $t\in\mathbb{N},Aid(X_{t})\geq 0$, this scenario will cease to occur after a finite times of updates.
Therefore, after updating precarious countries for finite time steps, all countries’ states will remain unchanged. Define
$\mathbf{S_{2}}=\Big\{\tilde{X}\in\mathbf{S_{1}}\Big|$ $\{X_{t}\}_{t\in\mathbb{N}}=\mathbf{X}(\tilde{X},\{i_{t}\}_{t\in\mathbb{N}^{+}})\text{ satisfies that }$
$sgn(s_{j}(X_{t}))=sgn(s_{j}(\tilde{X}))\text{ for all }t\in\mathbb{N}$
$\text{and }j\in\mathcal{V}\text{ if }s_{i_{t}}(X_{t-1})=0,\forall\ t\in\mathbb{N}^{+}.\Big\}.$
By definition, $\mathbf{S_{2}}$ is a subset of $\mathbf{S_{1}}$ with the following property: Given a strategy matrix in $\mathbf{S_{2}}$, if a precarious country is activated and updates its strategy, no country’s state will be changed. Moreover, there exist a sequence of strategy matrices $\{X^{(1)}_{t}\}_{0\leq t\leq T_{2}}=\mathbf{X}(X^{(1)},\{i_{t}\}_{1\leq t\leq T_{2}1})$ where $s_{i_{t}}(X^{(1)}_{t-1})\geq 0$ for any $t\leq T_{2}$ and $X^{(1)}_{T_{2}}\in\mathbf{S_{2}}$.
Hence, we have established that for any strategy matrix $X^{(0)}\in\mathbf{S_{1}}$, there exists a sequence of strategy matrices $\{X^{(0)}_{t}\}_{0\leq t\leq T_{1}+T_{2}}=\mathbf{X}(X^{(0)},\{i_{t}\}_{1\leq t\leq T_{1}+T_{2}})$ satisfying that $X^{(0)}_{T_{1}}\in\mathbf{S_{1}}$ and $X^{(0)}_{T_{1}+T_{2}}\in\mathbf{S_{2}}$. In the following two steps, we will prove two properties of the matrices in $\mathbf{S_{2}}$ to establish that, for any given matrix $X\in\mathbf{S_{2}}$, there exists an update sequence with finite activated countries that leads it to a stable state.
In step 2, we prove that there is no dangerous country given any strategy matrix in $\mathbf{S_{2}}$. Denote by $\mathbf{V}_{s}(X),\mathbf{V}_{p}(X),\text{ and }\mathbf{V}_{d}(X)$ the sets of safe countries, precarious countries, and dangerous countries respectively with the strategy matrix $X$. For any matrix $X^{(2)}\in\mathbf{S_{2}}$, assume that $\mathbf{V}_{p}(X^{(2)})=\{v^{p}_{1},v^{p}_{2},...,v^{p}_{n_{p}}\}$ and $\mathbf{V}_{s}(X^{(2)})=\{v^{s}_{1},v^{s}_{2},...,v^{s}_{n_{s}}\}$. We construct a sequence of strategy matrices $\{X^{(2)}_{t}\}_{0\leq t\leq n_{p}+n_{s}}=\mathbf{X}(X^{(2)},\{i_{t}\}_{1\leq t\leq n_{p}+n_{s}})\text{ where }i_{t}=v^{p}_{t},\text{ for all }1\leq t\leq n_{p}$ and $i_{t}=v^{s}_{t-n_{p}},\text{ for all }n_{p}+1\leq t\leq n_{p}+n_{s}$. We will prove that
$sgn(s_{j}(X^{(2)}_{t}))=sgn(s_{j}(X^{(2)}))$
for all $j\in\mathcal{V}$ and $0\leq t\leq n_{p}+n_{s}$. That is, no country’s state changes under the update sequence $\{i_{t}\}$ that we construct. According to the definition of $\mathbf{S}_{2}$, we have that $X^{(2)}_{1}\in\mathbf{S_{2}}$ and
$sgn(s_{j}(X^{(2)}_{1}))=sgn(s_{j}(X^{(2)}))$
for all $j\in\mathcal{V}$. Hence $s_{i_{2}}(X^{(2)}_{1})=0$. By iteration, we can obtain that, for any $j\in\mathcal{V}$ and $1\leq t\leq n_{p}$, $X^{(2)}_{t}\in\mathbf{S}_{2}$ and
$sgn(s_{j}(X^{(2)}_{t}))=sgn(s_{j}(X^{(2)}_{t-1}))=sgn(s_{j}(X^{(2)})).$
Since $\mathbf{S}_{2}\subseteq\mathbf{S}_{1}$, $X^{(2)}_{n_{s}}\in\mathbf{S}_{2}\subseteq\mathbf{S}_{1}$. Meanwhile, $sgn(s_{i_{n_{p}+1}}(X_{n_{p}})=sgn(s_{i_{n_{p}+1}}(X^{(2)}))=1$. According to the definition of $\mathbf{S}_{1}$, we have that $X_{n_{p}+1}\in H(X_{n_{p}},i_{n_{p}+1})$ belongs to $\mathbf{S}_{1}$ and
$N_{p}(\mathcal{V},X_{n_{p}+1})=N_{p}(\mathcal{V},X_{n_{p}}).$
Accoding to lemma 3.11,
$0\leq$ $u_{i_{n_{p}+1}}(X_{n_{p}+1})-u_{i_{n_{p}+1}}(X_{n_{p}})$
$\leq$ $N_{p}(\mathcal{V}\setminus\{i_{n_{p}+1}\},X_{n_{s}+1})$
$-N_{p}(\mathcal{V}\setminus\{i_{n_{p}+1}\},X_{n_{p}})$
$=$ $N_{p}(\mathcal{V},X_{n_{p}+1})-N_{p}(\mathcal{V},X_{n_{p}})=0$
Therefore, the update of country $i_{n_{p}+1}$’s strategy follow Rule 3. Moreover, it does not lead to a state change in any country. That is,
$sgn(s_{j}(X^{(2)}_{n_{p}+1}))$ $=sgn(s_{j}(X^{(2)}_{n_{p}}))$
$=sgn(s_{j}(X^{(2)}))$
for all $j\in\mathcal{V}$. Hence $s_{i_{n_{p}+2}}(X^{(2)}_{n_{p}+1})>0$. By iteration, we can obtain that, for any $j\in\mathcal{V}$ and $n_{p}+1\leq t\leq n_{p}+n_{s}$, $X^{(2)}_{t}\in\mathbf{S}_{1}$ and
$sgn(s_{j}(X^{(2)}_{t}))=sgn(s_{j}(X^{(2)}_{t-1}))=sgn(s_{j}(X^{(2)})).$
It follows from the above that, no state change is changed in any country under the chosen update sequence. This implies that in every step of the update, the utility of the updating country keep constant, so each update satisfies Rule 3. In addition to designing the update sequence $\{i_{t}\}$, we also construct a sequence of feasible strategies. For any $t\in\{1,2,\cdots,n_{p}+n_{s}\}$, the activated country reclaims all the power that is used to attack dangerous countries. As a result, we obtain a strategy matrix at time $n_{p}+n_{s}$ denoted as $X^{*}$, where no power is allocated by safe or precarious countries to attack dangerous countries, and all countries’ states keep unchanged during this process.
If $\mathbf{V}_{d}(X^{*})\neq\emptyset$, then:
$\forall i\in\mathbf{V}_{d}(X^{*}),\sum_{j\in\mathcal{F}_{i}}x^{*}_{ji}+\sum_{j\in\mathcal{A}_{i}}x^{*}_{ij}-\sum_{j\in\mathcal{A}_{i}}x^{*}_{ji}<0.$
Let the above formula be summed over all dangerous countries, then:
$\sum_{i\in\mathbf{V}_{d}(X^{*})}(\sum_{j\in\mathcal{F}_{i}}x^{*}_{ji}+\sum_{j\in\mathcal{A}_{i}}x^{*}_{ij}-\sum_{j\in\mathcal{A}_{i}}x^{*}_{ji})<0.$
Considering that the allocated power is non-negative and there is no power used to attack dangerous countries by safe or precarious countries, then:
$0$ $>\sum_{i\in\mathbf{V}_{d}(X^{*})}(\sum_{j\in\mathcal{F}_{i}}x^{*}_{ji}+\sum_{j\in\mathcal{A}_{i}}x^{*}_{ij}-\sum_{j\in\mathcal{A}_{i}}x^{*}_{ji})$
$\geq\sum_{i\in\mathbf{V}_{d}(X^{*})}\sum_{j\in\mathcal{A}_{i}}x^{*}_{ij}-\sum_{i\in\mathbf{V}_{d}}\sum_{j\in\mathcal{A}_{i}}x^{*}_{ji}$
$=\sum_{i\in\mathbf{V}_{d}(X^{*})}\sum_{j\in\mathcal{A}_{i}}x^{*}_{ij}-\sum_{i\in\mathbf{V}_{d}}\sum_{j\in\mathcal{A}_{i}\cap\mathbf{V}_{d}}x^{*}_{ji}\geq 0,$
which leads to contradiction. Therefore, $\mathbf{V}_{d}(X^{*})=\emptyset$, implying that there is no dangerous country given any strategy matrix $X^{(2)}\in\mathbf{S_{2}}$.
In step 3, we prove that, starting from any strategy matrix $X^{(3)}\in\mathbf{S_{2}}$, no country’s state changes regardless of the update sequence. Suppose that there exists an sequence of strategy matrices $\{X^{(3)}_{t}\}_{t\in\mathbb{N}}={\mathbf{X}}(X^{(3)},\{i_{t}\}_{t\in\mathbb{N}_{+}})$ along which at least one country’s state is changed. Let $t_{0}$ be the first time instance when the change of a country’s state happens. By definition of $\mathbf{S}_{1}$,
$X^{(3)}_{t}\in\mathbf{S}_{1},\text{ for all }0\leq t\leq t_{0}-1.$
Meanwhile, country $i_{t}$ must be in the precarious state at time $t$ and before, since updating a safe country does not change any country’s state and there is no dangerous country with the strategy matrix $X^{(3)}_{t_{0}-1}\in S_{1}$ . Given that no country’s state has changed before time $t_{0}$, we have
$p_{i}(t_{2})\geq p_{i}(t_{1}),\ \text{for any}\ i\in\mathcal{V}\ \text{and any}\ 0\leq t_{1}\leq t_{2}\leq t_{0},$
which means that no country’ power decreases, and
$p_{i_{t_{0}}}(t)=p_{i_{t_{0}}}(0),\forall\ 0\leq t\leq t_{0}-1,$
since country $i_{t_{0}}$ must be precarious. Since no country’s state change before time $t_{0}$, the updating country‘s utility keep constant at each time. Hence, its strategy update satisfies Rule 3, namely, it reclaims the power used to attack antagonistic countries. This implies that
$s_{i}(X^{(3)}_{t_{2}})\geq s_{i}(X^{(3)}_{t_{1}})\geq 0,$
for any $0\leq t_{1}\leq t_{2}\leq t_{0}-1$ and all $i\in\mathcal{V}$.
If there are changes in some countries’ states when the precarious country $i_{t_{0}}$ is selected to update at time $t_{0}$, then $u_{i_{t_{0}}}(X^{(3)}_{t})\geq u_{i_{t_{0}}}(X^{(3)}_{t-1})+1$, indicating that country $i_{t_{0}}$ reallocates power to strike other safe antagonistic countries in the update power-allocation strategy. We call the update power-allocation strategy used by country $i_{t_{0}}$ at time $t_{0}$ status-quo-breaking strategy and denote by $x_{t_{0}}$
Based on the matrix $X^{(3)}_{0}$, let $\hat{X}^{(3)}_{0}$ be the strategy matrix that only changes country $i_{t_{0}}$’s power-allocation strategy to the status-quo-breaking strategy $x_{t_{0}}$. Let us consider that how the country $i_{t_{0}}$’s utility changes if country $i_{t_{0}}$ choose the status-quo-breaking strategy before time $t_{0}$.
If $u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})\geq u_{i_{t_{0}}}(X^{(3)}_{0})+1$, there exists an available power-allocation strategy for country $i_{t_{0}}$ to strictly increase its utility when it is selected to update at time $1$, which contradicts the property of the set $\mathbf{S_{2}}$. Hence $u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})\leq u_{i_{t_{0}}}(X^{(3)}_{0})$. Let $u^{\triangle}=u_{i_{t_{0}}}(X^{(3)}_{0})-u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})$.
According to the Rule 3 of update strategy, any updating country only decreases its allocation power used to attack antagonistic countries and not change the allocation power used to support other friendly countries. Therefore,
$s_{j}(X^{(3)}_{t_{0}-1})+x^{(3)}_{i_{t_{0}}j}(t_{0}-1)\geq s_{j}(X^{(3)}_{0})+x^{(3)}_{i_{t_{0}}j}(0)$
with any $j\in\mathcal{A}_{i_{t_{0}}}$. Then,
$\sum_{j\in\mathcal{A}_{i_{t_{0}}}}u_{ji_{t_{0}}}(X^{(3)}_{t_{0}})\leq\sum_{j\in\mathcal{A}_{i_{t_{0}}}}u_{ji_{t_{0}}}(\hat{X}^{(3)}_{0})$.
Thus,
$\sum_{j\in\mathcal{F}_{i_{t_{0}}}}u_{ji_{t_{0}}}(X^{(3)}_{t_{0}})$= $\ u_{i_{t_{0}}}(X^{(3)}_{t_{0}})-\sum_{j\in\mathcal{A}_{i_{t_{0}}}}u_{ji_{t_{0}}}(X^{(3)}_{t_{0}})$
$\geq$ $\ u_{i_{t_{0}}}(X^{(3)}_{0})+1-\sum_{j\in\mathcal{A}_{i_{t_{0}}}}u_{ji_{t_{0}}}(\hat{X}^{(3)}_{0})$
$=$ $\sum_{j\in\mathcal{F}_{i_{t_{0}}}}u_{ji_{t_{0}}}(\hat{X}^{(3)}_{0})+u^{\triangle}+1.$
So the reason that country $i_{t_{0}}$ can not increase its utility by applying the status-quo-breaking strategy at time 0 is that at least $u^{\triangle}+1$ country $i_{t_{0}}$’s friendly countries transfers from safety to danger if country $i_{t_{0}}$ takes the status-quo-breaking strategy at time 0, but not being dangerous given the strategy matrix $X^{(3)}_{t_{0}}$. Denote by $i_{f_{1}},i_{f_{2}},...,i_{f_{u^{\triangle}+1}}$ this kind of country $i_{t_{0}}$’s friendly countries. Let $t^{*}_{k}:=x^{(3)}_{i_{t_{0}}i_{f_{k}}}(0)-x^{(3)}_{i_{t_{0}}i_{f_{k}}}({t_{0}})-s_{i_{f_{k}}}(X^{(3)}_{0})$ and $t^{*}=\sum_{k=1}^{u^{\triangle}+1}t^{*}_{k}+1$. We can construct a sequence of strategy matrices $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}}=\mathbf{X}(X^{(3)},\{i_{t}\}_{1\leq t\leq t^{*}})$ which we denote by $\bar{\mathbf{X}}^{(3)}$. Let $i_{t}=i_{f_{k}}$ for any $\sum_{j=1}^{k-1}t^{*}_{j}\leq t\leq\sum_{j=1}^{k}t^{*}_{j}-1$ and all $1\leq k\leq u^{\triangle}+1$. Let $i_{t^{*}}=i_{t_{0}}$. Meanwhile, before time $t^{*}$, let the updating country’s power increase by $1$ and make the updating country to adopt a feasible strategy in which the allocation of power to the other countries remains unchanged, and any additional power is solely used to protect itself. Let $i_{t_{0}}$ choose the status-quo-breaking strategy at time $t^{*}$. Then for all $1\leq k\leq u^{\triangle}+1$, country $i_{f_{k}}$’s state will transfer from safe to precarious during the update at time $t^{*}$. Since the updating countries befor $t^{*}$ are all in the safe state, according to what was explained in step 2, no country’s state will change before $t^{*}$ with the sequence of strategy matrices $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}}$, which means that $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}-1}$ is a feasible sequence. Notice that
$u_{i_{t_{0}}}(\bar{X}^{(3)}_{t_{*}})=\sum_{j\in\mathcal{A}_{i_{t_{0}}}}u_{ji_{t_{0}}}(\bar{X}^{(3)}_{t_{*}})+\sum_{j\in\mathcal{F}_{i_{t_{0}}}}u_{ji_{t_{0}}}(\bar{X}^{(3)}_{t_{*}})$
$\geq\sum_{j\in\mathcal{A}_{i_{t_{0}}}}u_{ji_{t_{0}}}(\bar{X}^{(3)}_{0})+\sum_{j\in\mathcal{F}_{i_{t_{0}}}}u_{ji_{t_{0}}}(\bar{X}^{(3)}_{0})+u^{\triangle}+1$
$=u_{i_{t_{0}}}(\bar{X}^{(3)}_{0})+1=u_{i_{t_{0}}}(\bar{X}^{(3)}_{t^{*}-1})+1.$
Hence the status-quo-breaking strategy is a feasible strategy at time $t^{*}$ for country $i_{t_{0}}$.
According to the proof in Lemma 3.11:
$N_{p}(\mathbf{C}^{-}_{i_{t_{0}},{t^{*}}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}})$- $N_{p}(\mathbf{C}^{-}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}-1})$
$\geq-card(\mathbf{C}^{-}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)})),$
$N_{p}(\mathbf{C}^{+}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}})$- $N_{p}(\mathbf{C}^{+}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}-1})$
$\geq card(\mathbf{C}^{+}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)})),$
$N_{p}(\mathbf{C}^{0}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}})$- $N_{p}(\mathbf{C}^{0}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}-1})\geq 1.$
The right side of the third equation is 1 instead of 0 because at least one of country $i_{t_{0}}$’s friendly country is in the safe state at time $t^{*}-1$ and in the precarious state at time $t^{*}$.
Therefore:
$N_{p}(\mathcal{V},t^{*})-N_{p}(\mathcal{V},t^{*}-1)$
$=N_{p}(\mathcal{V}\setminus\{i_{i_{0}}\},t^{*})-N_{p}(\mathcal{V}\setminus\{i_{i_{0}}\},t^{*}-1)-1$
$=N_{p}(\mathbf{C}^{-}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}})-N_{p}(\mathbf{C}^{-}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}-1})$
$+N_{p}(\mathbf{C}^{+}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}})-N_{p}(\mathbf{C}^{+}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}-1})$
$+N_{p}(\mathbf{C}^{0}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}})-N_{p}(\mathbf{C}^{0}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}),X_{t^{*}-1})$
$-1$
$\geq card(\mathbf{C}^{+}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}))-card(\mathbf{C}^{-}_{i_{t_{0}},t^{*}}(\bar{\mathbf{X}}^{(3)}))+1-1$
$=u_{i_{t_{0}}}(X_{t^{*}})-u_{i_{t_{0}}}(X_{t^{*}-1})+1-1$
$\geq 1.$
which contradicts the property of the set $\mathbf{S_{1}}$.
Now, we can assert that given any strategy matrix $X\in\mathbf{S_{2}}$, no country’s state changes no matter what the following update sequence is. Since all countries are not dangerous, the power of all countries will continue to increase. Moreover, the power of countries will converge, which is followed by the fact that the countries’ power does not decrease and can not exceed the upper bound $K$. Considering that the countries’ power will converge in finite times and the state of all countries keep constant, the strategy matrix will converge in finite times, too. Meanwhile, the convergence matrix is a Nash Equilibrium.
In conclusion, we have established that, given any initial strategy matrix, there exists a sequence of finite update countries such that the initial matrix will be transitioned to a steady state. According to Lemma 3.10, the dynamic process will almost surely converge to a steady state. Furthermore, the resulting steady-state strategy matrix constitutes a Nash Equilibrium of the power-allocation game defined in Section 2.1.
4 Numerical and Empirical Studies
4.1 Simulation of the static model: which countries will survive?
Regarding the static power-allocation game introduced in Section 2.1, although Theorem 1 establishes the existence of a pure-strategy Nash equilibrium, the game in fact could admit multiple Nash equilibria. Intuitively, a country’s states at all these Nash equilibria reflect its overall safety level. In this subsection, we approximately compute the countries’ safety levels via the following simulations: For each simulation, at each time step, a country is uniformly randomly picked to update its power allocation as the best response that maximizes the following utility function:
$u_{i}(X)=\begin{cases}0,&\text{if}~s_{i}(X)<0,\\ \sum_{j\in\mathcal{F}_{i}\cup\mathcal{A}_{i}}u_{ji}(X),&\text{if}~s_{i}(X)\geq 0.\end{cases}$
The simulation terminates after $10^{7}$ rounds of iterations and the final state is an approximation of one Nash equilibria. We run such simulations for 10000 times. Based on the simulation results, a country’s likelihood of survival is numerically defined as the proportion of the finals states where its state is NOT dangerous.
We test our model’s predictive power via historic data. The signed networks are constructed based on the data collected from the Correlates of War project [MaozEtAl2019], which records countries’ conflicts and cooperative relationships in every year. Two countries are considered friendly if the cooperative incidents between them outscore the conflicts, and vice versa. The signed networks of international relations in each year from 1939 to 1944, i.e., during the Second World War, are generated, see Fig.1A for the year of 1940 as an example. Countries’ powers in each of these years are generated according to the composite index for national capability (CINC) dataset [singer2012capability]. The distributions of all the countries’ likelihoods of survival from 1939 to 1944, computed via the approach mentioned in the previous paragraph, are shown in Fig.1B. Based on whether a country’s likelihood of survival exceeds 0.5, we predict whether battles of the Second World War will occur within its territory in the following year.
Figure 1: A) is the graph of international relationships in 1940, where the green edges represent that the two countries are friendly countries and the red edges represent that they are antagonistic. B)- G) represent respectively the distribution histogram of survival likelihood from 1939 to 1944.
While the above prediction accuracies are not impressively high, our model, considering its simplicity, serves as a good benchmark and a promising starting point for further improvement. Moreover, the prediction accuracies are somehow robust to the choice of the threshold value (0.5 as mentioned above), since the simulation results exhibits a desriable feature that the computed likelihoods of survival are not ambiguous but concentrate around the values 0 or 1. The only exception is the year of 1941. In that year, the likelihoods of survival for some countries clustered around 0.5, a feature consistent with the historical record: The overall course of the war became more uncertain, as the United States had formally entered the war in 1941 but had not yet fully demonstrated the potential of its war industry.
4.2 Empirical support for the dynamic model of power-strategy co-evolution
The dynamic model of the co-evolution between countries’ powers and power allocations, proposed and studied in Section 3, is based on a critical assumption that a country’s power increases when it is safe, and vice versa. In this subsection, we provide empirical evidence that support this assumption. Based on real-world historic data, we test whether a country’s likelihood of survival predicts an increase or decrease of its power. Similar to what we have done in Section 4.1, signed international-relation networks for each year from 1955 to 2002 are generated based on the Correlates of War dataset [MaozEtAl2019]. Every country’s power in each year is quantized based on the Penn World Table (PWT) dataset (version 10.0) [feenstraNextGenerationPenn2015].
Figure 2: A) is the heat map of $p_{avg}$ and $(q_{e},q_{n})$, where the color of each point represents the number of average power. B)- E) represent scatter plots of the average power and the frustration of network given $(q_{e},q_{n})$, where $(q_{e},q_{n})=(0.2,0.2)$ in B), $(q_{e},q_{n})=(0.2,0.8)$ in C), $(q_{e},q_{n})=(0.8,0.2)$ in D), and $(q_{e},q_{n})=(0.8,0.8)$ in E). F) is the heat map of Gini coefficient and $(q_{e},q_{n})$, where the color of each point represents the number of Gini coefficient. G)- J) represent scatter plots of the average power and the frustration of network given $(q_{e},q_{n})$, where $(q_{e},q_{n})=(0.2,0.2)$ in G), $(q_{e},q_{n})=(0.2,0.8)$ in H), $(q_{e},q_{n})=(0.8,0.2)$ in I), and $(q_{e},q_{n})=(0.8,0.8)$ in J)
Now we aim to determine which kind of signed graphs make countries thrive. The algorithm used here is based on the dynamics defined in Section 3.1. In a simulation, we randomly generate a relational network to represent countries’ relationships. Then we randomly initialize countries’ power and a strategy matrix as a starting point. At each time step, we randomly choose a country to update its power and strategy according to the update rules defined in Section 3.1. Finally, we record all countries’ power to calculate its mean and Gini coefficient. After a pre-set number of simulation rounds, we calculate the mean of the means and gini coefficient of all countries’ power across all simulations and call them the average power and gini coefficient respectively. We construct networks by connecting any two nodes with probability $q_{e}$ and then assigning a negative sign to each edge with probability $q_{n}$.
The simulation results are represented in Figure  2. From the figure, we can see that the ratio of negative edges will decrease the average power and when the ratio of negative edges remains constant, the average power will decrease when the number of edges increases, indicating that the influence of the negative edges on the average power is more significant than the influence of the same ratio of positive edges. Meanwhile, the Gini coefficient is affected in the exact opposite way to the average coefficient. Given the $q_{e}$ and $q_{n}$, the average power and frustration are negatively correlated, which means that the more balanced the graph structure, the greater the average power. When $q_{n}=0.2$, the gini coefficient and frustration are positively correlated; when $q_{e}=0.2,q_{n}=0.8$, they are negatively correlated.
5 Conclusion and Further Discussion
Similar to the approach in[LiMorse2022], the first part of this paper establishes the existence of a pure strategy Nash equilibrium for a new type of power allocation games. Instead of relying on the classic theorems, such as fix-point theorems and the theorems which are applied to matrix games, we employ an algorithm that directly constructs such an equilibrium for the games. On this basis, the second part of the paper generalizes the model to a dynamic process and establishes that it almost surely converges. Meanwhile, the convergence matrix coincides with the Nash equilibrium of the corresponding static model under identical conditions. Additionally, we use real-world data to simulate predictions for countries’ survivability. We also perform simulations related to the dynamic model to predict whether national GDPs will increase or decrease. Lastly, we perform simulations on the dynamic system to figure out which signed network derived from historical international relations can lead to the prosperity of all countries.
There are several promising avenues for future research. One possible extension of the model is to assign different weights to the relationships between countries, which may lead to the emergence of more diverse results. Meanwhile, inverse inference and intention recognition under incomplete information can also be incorporated into our model for further analysis.